### v1

# Permutation wordle

Abstract.
We introduce a guessing game, permutation Wordle, in which a guesser attempts to recover a hidden permutation in $S_{n}$. In each round, the guesser guesses a permutation (using information from previous rounds) and is told which entries of that permutation are correct. We describe a natural guessing strategy, which we believe to be optimal. We show that the number of permutations this strategy solves in $k+1$ rounds is the Eulerian number $A(n,k)$.
We also describe an extension to suited permutations: the setter chooses a permutation in $S_{n}$ and also a coloring of $[n]$ using $s$ colors. We generalize our strategy, give a recurrence for the number of suited permutations solved in $k+1$ rounds, and relate these numbers to the Eulerian numbers. In the case of two suits, or signed permutations, we also relate these numbers to the Eulerian numbers of type B.
1. Introduction
Consider the following game: A setter chooses a permutation on $n$ objects; say, cards labeled $1$ through $n$. Your job is to guess this hidden permutation. To make a guess, you arrange the $n$ cards, and the setter tells you which ones are in the correct place; you win if all $n$ are correct. What is the optimal strategy? How many guesses do you expect to need?
We call this game “permutation Wordle”, because the guesser’s feedback is similar to that in Josh Wardle’s popular online game Wordle. (For connections to other games, and to prior analysis of those games, see Section 2.) We propose a simple strategy, which we call “CircularShift”: After each guess, leave in place the cards known to be correct. Shift each incorrect card one space to the right (skipping over the correct cards), cycling the rightmost incorrect card around to the space formerly occupied by the leftmost. We prove the following:
Theorem 1.1.
The number of permutations solved in $k+1$ guesses by CircularShift is the Eulerian number $A(n,k)$: the number of permutations in $S_{n}$ with $k$ exceedances.
The Eulerian numbers come up in a variety of contexts in combinatorics; we discuss several definitions, and prove Theorem 1.1, in Section 3.
We conjecture that CircularShift is optimal in the following sense: Suppose the setter chooses a permutation uniformly from $S_{n}$. Given a strategy ${\mathcal{S}}$, for each $r$, let $P_{r}({\mathcal{S}})$ denote the probability that ${\mathcal{S}}$ guesses the correct permutation within $r$ guesses.
Conjecture 1.2.
$P_{r}(\textnormal{{CircularShift}})$ attains the maximal value of $P_{r}$ for all $r\geq 1$.
We also consider a game we call “suited permutation Wordle”; imagine that there is a deck of $ns$ cards labeled $1$ through $n$ in each of $s$ different suits, and that the setter has chosen an arrangement of cards in which each number appears once. Your guess is also an arrangement of $n$ cards in which each number appears once, and, again, you learn which cards are correct. The game is harder because some of your guessed cards may not be in the setter’s arrangement at all; you may have the wrong suit. In Section 4 we describe and analyze a generalization of the CircularShift strategy, and we show that Eulerian numbers continue to play a role.
In the case $s=2$, these are more commonly called signed permutations. In this particular case, if we look at the number of permutations solved in $12k-$ 1 or $2k$ rounds, we obtain a multiple of the Eulerian numbers. However, if we look at the number solved in $2k$ or $12k+$ 1 rounds, we obtain the Eulerian numbers of type B. This game thus provides a surprising connection between the two sequences.
2. Guessing games
In the classic game Mastermind, invented by Mordecai Meirowitz, the setter chooses a “code”: a string of length $n$ from an alphabet of size $m$ (which we identify as $[m]=\{1,\ldots,m\}$). In each round, the guesser tries a string in $[m]^{n}$, and the setter responds with two numbers: the number of entries that are correct (right number, right place) and the number of additional entries that are present elsewhere in the code (right number, wrong place). Donald Knuth [Knu77] showed that, for the original game ($n=4$ and $m=6$), five rounds are sufficient for the guesser to guess the code. Many others have evaluated the game for various values of $n$ and $m$. The typical emphasis is on a worst-case analysis; one can imagine the setter adversarially altering the code to be consistent with all information presented thus far, with the goal of maximizing the number of rounds.
One Mastermind variant is “permutation Mastermind”, in which $m=n$ and both the setter’s code and the guesser’s guesses must contain each value exactly once. (In this variant, the setter simply responds with the number $c$ of correct entries, since the number in the wrong place is simply $n-c$.) El Ouali et al. [EOGSS18] have shown that $n\log_{2}n+O(n)$ queries suffice.
In 2020, a new online game became popular: Wordle, by Josh Wardle. The setup is the same as Mastermind, but the feedback is different: the setter indicates which entries are correct and in the correct place, and which additional entries are present in the code elswhere. [^1] The rules for handling cases where a value occurs multiple times in the code or the guess are irrelevant to our permutation problem. In the original Wordle, the alphabet is the letters A–Z, and the code and guesses are required to be 5-letter words in English. There are numerous other variants, including a recent one called “OrderUp” [Wag24] in which the code and guesses are 7-long permutations.
Li and Zhu [LZ24] have analyzed a generalized version of Wordle, with a code from the alphabet $[m]^{n}$. Among other things, they show that, if $m=n$ and the code is required to use each value exactly once, then the guesser needs $n$ guesses in the worst case; this is true even if the guesses are are allowed to have values multiple times. They give a simple strategy with a matching upper bound of $n$ guesses for the general case $m=n$: the first $n-1$ guesses are the strings $1^{n}$, $2^{n}$, $\ldots$, $(n-1)^{n}$, at which point the guesser knows the code. We will show that, in fact, the guesser can guess an $n$-long permutation in at most $n$ guesses even when the guesses are required to be permutations.
As with the Mastermind analyses, Li and Zhu emphasize worst-case behavior; any average-case analysis would need to consider a probability distribution over codes. As we will see below, in the case of permutations, all codes are of equal difficulty; we may as well assume the setter chooses a random code. Therefore, we consider average-case behavior; or, more generally, for each $r\leq n$, we consider the probability that the guesser succeeds within $r$ guesses.
3. Eulerian numbers and CircularShift
We will describe a simple strategy, CircularShift, which solves permutation Wordle in at most $n$ guesses. As noted above, Li and Zhu [LZ24] have previously shown that $n$ guesses are necessary. In our setting, we can investigate average-case behavior.
Recall that the setter chooses a permutation $\sigma\in S_{n}$; let $\{\gamma_{r}\}$ for $r\geq 1$ denote the guesser’s permutations. The feedback in round $r$ is the set $\{i\mid\gamma_{r}(i)=\sigma(i)\}$, or, equivalently, the set of fixed points of $\sigma\gamma_{r}^{-1}$. If the setter had instead chosen $\sigma\tau$, and the guesser’s $r$th guess were $\gamma_{r}\tau$, the feedback sequence would be exactly the same. Equivalently, the guesser could choose a uniform random $\tau$, and multiply each guess by $\tau$ throughout. We can therefore assume that the setter chooses $\sigma$ uniformly at random, and, for simplicity, we let $\gamma_{1}$ be the identity permutation.
We discuss average-case behavior; we compute the probability of solving permutation Wordle in exactly $r$ guesses for various values of $r$. For example, consider the case $n=3$. Assume the setter’s $\sigma$ is chosen randomly, and that our first guess $\gamma_{1}$ is the identity permutation $(1,2,3)$; we then learn the set of fixed points of $\sigma$. If $\sigma$ is the identity, we are immediately done; if $\sigma$ has one fixed point, we will be able to guess it in round 2. If $\sigma$ is one of the two 3-cycles, then our best strategy is to guess a 3-cycle in round 2; we will either be correct, or will need a third guess. Overall, the probability of finding $\sigma$ in one round is $1/6$; in two rounds, $4/6$; in three rounds, $1/6$.
When $n=4$, the analysis becomes slightly more complicated. Again, let $\gamma_{1}$ be the identity. If $\sigma$ is the identity, we finish in one round; if it has two fixed points, we need two rounds; if it has one fixed point, we need two or three rounds. When $\sigma$ is one of the nine derangements, the optimal strategy turns out to be to let $\gamma_{2}$ be a $4$-cycle. If $\sigma=\gamma_{2}$, we are done in two rounds; if $\sigma\gamma_{2}^{-1}$ has any fixed points, we can deduce $\sigma$ in the third round; otherwise, $\sigma\in\{\gamma_{2}^{2},\gamma_{2}^{3}\}$, and we need three or four rounds.
Overall, for $n=4$, the probability of finishing in one round is $1/24$; in two rounds, $11/24$; in three rounds, $11/24$; in four rounds, $1/24$. Equivalently, the probabilities $P_{r}$ of finishing it at most $r$ rounds are $1/24$, $1/2$, $23/24$, $1$. One can show that for $n=4$ this strategy dominates all others in the following sense: if we compute the probabilities $P^{\prime}_{r}$ for any other strategy, then $P_{r}\geq P^{\prime}_{r}$ for all $r$.
3.1. The strategy CircularShift
We now extend this strategy to larger $n$. We describe a deterministic approach to generating the guesses $\gamma_{r}$; as noted above, for a randomized version, we could choose a permutation $\tau$ and guess $\gamma^{\prime}_{r}=\gamma_{r}\tau$.
Definition 3.1 (CircularShift).
Let $\gamma_{1}=(1,2,\ldots,n)$. Suppose that, in round $r$, we submit guess $\gamma_{r}$ and receive feedback $U_{r}$, the set of values on which $\gamma_{r}$ agrees with the setter’s permutation $\sigma$. If $U_{r}=[n]$, we are done in $r$ rounds. Otherwise, we define $\gamma_{r+1}$ as follows: first, write $[n]\setminus U_{r}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$. Then
$\gamma_{r+1}(i):=\begin{cases}\gamma_{r}(i)&\text{if $i\in U_{r}$,}\\ \gamma_{r}(a_{j-1})&\text{if $i=a_{j}$ for some $j>1$,}\\ \gamma_{r}(a_{\ell})&\text{if $i=a_{1}$.}\end{cases}$
We will let $D(n,r)$ denote the number of permutations for which CircularShift succeeds in exactly $r$ rounds.
1 2 3 4 5 6 7 8 9
$\gamma_{1}$ 1 2 3 4 5 6 7 8 9
$\gamma_{2}$ 8 2 1 3 5 4 6 7 9
$\gamma_{3}$ 7 2 8 1 5 3 4 6 9
$\gamma_{4}$ 7 2 4 8 5 3 1 6 9
Table 1. The deterministic strategy permutation Wordle guessing the permutation $\sigma=(7,2,4,8,5,3,1,6,9)$. Correct values are shown in boldface; the underlines will be discussed in Section 3.3. The strategy succees in four rounds.
We work out an example in Table 1. In the first round, we guess the identity permutation and get three values correct. In round 2, we cyclically shift the remaining six values; none of them are correct, so we shift them again in round 3. At this point, we have six correct, and in round 4 we fix the remaining values.
The probability of succeeding in at most $r$ rounds is $\sum_{k=1}^{r}D(n,k)/n!$. The claim of Conjecture 1.2 is that these probabilities are all larger for CircularShift then for any other strategy. We have only proved this for small $n$ (specifically, $n\leq 4$). Computer experiments for $n=5,6$ have failed to find any better strategies.
We should note that there is a dual formulation of our problem: in this scenario, we imagine that we were able to choose $\tau=\sigma^{-1}$. That is, we start with a specified permutation, we perform the CircularShift operation on it $r$ times, and we end up with the identity. This dual approach is further removed from the original guessing game, but may make the analysis of CircularShift more natural for some readers. The two approaches are equivalent.
3.2. Eulerian numbers
The Eulerian numbers occur frequently in the combinatorics of permutations. One common definition involves counting the number of descents in a permutation: that is, the number of values $i$ for which $\pi(i)>\pi(i+1)$. For our purposes, it will be more convenient to use a formulation in terms of exceedances.
Definition 3.2.
For $\pi\in S_{n}$, a exceedance is a value $i$ for which $\pi(i)>i$. We let $\operatorname{Exc}(\pi)$ denote the set of exceedances: $\operatorname{Exc}(\pi):=\{i\mid\pi(i)>i\}$. We write $\operatorname{exc}(\pi)=\#\operatorname{Exc}(\pi)$.
For $0\leq k<n$, the Eulerian number $A(n,k)$ is #$\#\{\pi\in S_{n}\mid\operatorname{exc}(\pi)=k\}$.
For example, when $n=3$, $A(3,0)=1$ (the only such permutation is $(1,2,3)$) and $A(3,2)=1$ (the only such permutation is $(2,3,1)$); we then have $A(3,1)=4$.
Proposition 3.3.
The counts $A(n,k)$ satisfy the recurrence
(1) $A(n,k)=(k+1)A(n-1,k)+(n-k)A(n-1,k-1)$
with initial conditions $A(0,0)=1$ and $A(0,k)=0$ for $k\neq 0$.
Proof.
We induct on $n$; the unique element of $S_{0}$ has 00 exceedances, verifying the initial condition.
Every permutation $\pi\in S_{n}$ can be uniquely decomposed in the form $\alpha\beta$, where $\beta$ fixes $n$ and $\alpha$ is either the identity, or swaps $n$ with some value $j$. We view $\beta$ as an element of $S_{n-1}$.
How can $\pi$ have exactly $k$ exceedances? There are two possibilities. First, $\beta$ could have $k$ exceedances, and $\alpha$ is either the identity, or swaps $n$ with one of the $k$ places where $\beta$ already has an exceedance. Second, $\beta$ could have $k-1$ exceedances, and $\alpha$ swaps $n$ with one of the $n-k$ places where $\beta$ did not have an exceedance. This corresponds exactly to the desired reucrrence. ∎
It is straightforward to show that, if we had defined $A(n,k)$ in terms of the number of permutations with $k$ descents, the same recurrence would apply. Alternatively, Foata and Schützenberger [FS70] givee an explicit bijection between permutations with $k$ exceedances and those with $k$ descents.
Note that the recurrence implies the reflective symmetry $A(n,k)=A(n,n-k-1)$.
The Eulerian numbers show up in a variety of other ways; for example, if $Y_{i}$ are uniform random variables in the interval $(0,1)$, and we define $X=\left\lfloor{\sum_{i=1}^{n}Y_{i}}\right\rfloor$, then the probability that $X=k$ is given by $A(n,k)/n!$. We now show that these numbers also arise in permutation Wordle.
3.3. The main theorem
We now connect the performance of CircularShift with exceedances.
Theorem 3.4.
The number of guesses it takes CircularShift to guess a permutation $\sigma$ is $1+\operatorname{exc}(\sigma)$.
Note that Theorem 1.1 follows immediately from Theorem 3.4 and from Definition 3.2.
Proof of Theorem 3.4.
Let $\sigma$ be the setter’s permutation; we run CircularShift, producing a series of guesses $\gamma_{1}$, …, $\gamma_{r}$ for some $r$ with $\gamma_{r}=\sigma$. Define the set $R_{k}$ to be the values that $\gamma_{k}$ places further right than $\sigma$ does:
$R_{k}=\{i\mid\gamma_{k}^{-1}(i)>\sigma^{-1}(i)\}.$
(The numbers in $R_{k}$ are underlined in Table 1.) We again let $U_{k}$ denote the values that are already fixed; that is, those $i$ with $\gamma_{k}^{-1}(i)=\sigma^{-1}(i)$.
How does $R_{k}$ change as $k$ increases? Let $\ell$ denote the rightmost entry in $\gamma_{k}$ not in $U_{k}$; then $\ell$ is to the right of its position in $\sigma$, so $\ell\in R_{k}$. In $\gamma_{k+1}$, this value $\ell$ will cycle around to the first available position (that is, anything to its left is in $U_{k}$), so $\ell\not\in R_{k+1}$.
Any other values in $R_{k}$ will move further right, and will remain in $R_{k+1}$. Any values in $U_{k}$ will remain where they were. Finally, any other values (those $i$ with $\gamma_{k}^{-1}(i)>\sigma^{-1}(i)$) will move right one step, but cannot move past their location in $\sigma$, so they will not be in $R_{k+1}$. We conclude that $R_{k+1}=R_{k}\setminus\{\ell\}$, so #$\#R_{k+1}=\#R_{k}-1$.
Since $R_{r}$ is empty, the number of rounds used by CircularShift is $1+\#R_{1}$. Finally, $i\in R_{1}$ if and only if $i>\sigma^{-1}(i)$, or $\sigma^{-1}(i)\in\operatorname{Exc}(\sigma)$. ∎
Corollary 3.5.
The expected number of rounds for CircularShift is $(n+1)/2$.
Proof.
This follows immediately from the symmetry of the Eulerian numbers: $A(n,k)=A(n,n-k-1)$. ∎
This leads us to a weaker version of Conjecture 1.2:
Conjecture 3.6.
For any strategy for permutation Wordle, the average number of rounds is at least $(n+1)/2$.
4. Suited permutation Wordle
We now turn to the “suited” version of permutation Wordle. The setter’s $\sigma$, and the guesses $\gamma_{k}$, are each now a permutation in $S_{n}$ together with a suit assignment (or coloring) $[n]\to\{0,\ldots,s-1\}$. Equivalently, a suited permutation is a list of $n$ ordered pairs $((c_{1},i_{1}),\ldots,(c_{n},i_{n}))$, where $(i_{1},\ldots,i_{n})$ is a permutation and $0\leq c_{j}<s$. The feedback [^2] One interpretation of the Wordle rules would be that a suited number, or “card”, is colored green if it is the right place, and yellow if it occurs in the permutation but in a different place, but this leads to a less interesting math problem. will be as before: we are told the list of $j$ where our guess agrees with $\sigma$.
[…]
As with permutation Wordle, we can assume that the setter chooses $\sigma$ uniformly at random, and that our first guess $\gamma_{1}$ is the identity with all suits chosen to be 00.
Definition 4.1 (CircularShift for suited permutation Wordle).
Let $\gamma_{1}=((0,1),(0,2),\ldots,(0,n)$. Suppose that, in round $r$, we submit guess $\gamma_{r}$ and receive feedback $U_{r}$, the set of values on which $\gamma_{r}$ agrees with the setter’s permutation $\sigma$. If $U_{r}=[n]$, we are done in $r$ rounds. Otherwise, we define $\gamma_{r+1}$ as follows: first, write $[n]\setminus U_{r}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$. Write $\gamma_{r}(a_{1})=(c_{1},i_{1})$. Then
$\gamma_{r+1}(i):=\begin{cases}\gamma_{r}(i)&\text{if $i\in U_{r}$,}\\ \gamma_{r}(a_{j-1})&\text{if $i=a_{j}$ for some $j>1$,}\\ (c_{\ell}\mathbin{\boxminus}1,i_{\ell})&\text{if $i=a_{1}$.}\end{cases}$
We let $D_{s}(n,r)$ denote the number of permutations for which CircularShift succeeds in exactly $r+1$ rounds.
Conjecture 4.2.
For each $r$, the probability that CircularShift guess the correct suited permutation within $r$ guesses is optimal, that is, it is at least as large as the corresponding probability for any other strategy.
Our main result on suited permutation Wordle is a recurrence relation that can be used to compute $D_{s}(n,r)$:
Theorem 4.3.
The counts $D_{s}(n,r)$ satisfy the recurrence
(2) $D_{s}(n,r)=(r+1)D_{s}(n-1,r)+(sn-r)D_{s}(n-1,r-s)+\sum_{t=1}^{s-1}D_{s}(n-1,r-t)$
with initial conditions $D_{s}(0,0)=1$ and $D_{s}(0,r)=0$ for $r\neq 0$.
Note that, when $s=1$, this is exactly the recurrence for the Eulerian numbers from Proposition 3.3.
We first define a mapping $\phi$ from suited $n$-perms to non-suited $sn$-perms. We could explicitly work with integers in $[sn]$ rather than with ordered pairs $(c,i)$, with the mapping $(c,i)\mapsto i+cn$. However, it will be more convenient to work directly with ordered pairs, sorted lexicographically.
Definition 4.4.
Given an $s$-suited $n$-permutation $\tau=((c_{1},i_{1}),\ldots,(c_{n},i_{n}))$, we define the permutation $\phi(\tau)$ on the $sn$ possible ordered pairs by
$\phi(\tau)\colon(c,j)\mapsto(c_{j}\mathbin{\boxplus}c,i_{j}).$
We let $\Phi\subset S_{sn}$ denote the image of the set of all suited permutations under $\phi$.
Proposition 4.5.
If $\gamma_{1}$, …, $\gamma_{r}$ is the sequence of guesses for the suited permutation $\sigma$ following CircularShift, then $\phi(\gamma_{1}),\ldots,\phi(\gamma_{r})$ is the sequence of guesses for $\phi(\sigma)$ following CircularShift.
Proof.
This follows from the definition of the suited and non-suited versions of CircularShift. ∎
Corollary 4.6.
The count $D_{s}(n,r)$ is the number of permutations in $\Phi$ with exactly $r$ exceedances.
Lemma 4.7.
Let $\tau$ be a suited permutation $((c_{1},i_{1}),\ldots,(c_{n},i_{n}))$. For any $j$, let $J=\{(c,j)\mid 0\leq c\leq s-1\}$. Let $E$ denote $\operatorname{Exc}(\phi(\tau))\cap J$, the exceedances of $\phi(\tau)$ with indices in the set $J$. Then $E=\{(c,j)\mid 0\leq c<\ell\}$, where
=$\ell=\begin{cases}0&\text{$c_{j}=0$ and $i_{j}\leq j$,}\\ s&\text{$c_{j}=0$ and $i_{j}>j$,}\\ s-c_{j}&\text{$c_{j}>0$}.\end{cases}$
In particular, #$\#E\equiv s-c_{j}\pmod{s}$.
Proof.
When $c_{j}=0$, then $\phi(\tau)$ maps $(c,j)\mapsto(c,i_{j})$. Hence, either all $s$ values are exceedances (when $i_{j}>j$) or none of them are.
When $c_{j}>0$, then, regardless of $i_{j}$, $(c_{j}+c,i_{j})$ is lexicographically larger than $(c,j)$. So $(0,j)\in\operatorname{Exc}(\phi(\tau))$, and $(c,j)\in\operatorname{Exc}(\phi(\tau))$ as long as $c+c_{j}\leq s-1$, or $c<s-c_{j}$. When $c\geq s-c_{j}$, then $(c_{j}+c-s,i_{j})$ is lexicographically smaller than $(c,j)$, so no additional values of $c$ are exceedances. ∎
We are now ready to prove the recurrence for $D_{s}(n,r)$.
Proof of Theorem 4.3.
We generalize the argument of Proposition 3.3. We induct on $n$; the case $n=0$ is clear.
Every suited $n$-permutation $\eta$ can be generated from a suited $(n-1)$-permutation $\beta$. To begin, form $\eta$ from $\beta$ by appending $(c_{n},n)$ for some choice of $c_{n}$. Then, we have two cases: first, we could leave $(c_{n},n)$ where it is. Second, we place $(c_{n},n)$ in the $j$th position for some $j<n$, and we then place $(c^{\prime}_{j},i_{j})$ in the $n$th position, where $c^{\prime}_{j}=c_{j}\mathbin{\boxminus}c$. (The reason for this specific choice of $c^{\prime}_{j}$ will become appparent.) This process generates each suited $n$-permutation excatly once. Our goal is to count the number of ways to generate $\eta$ so that $\phi(\eta)$ has exactly $r$ exceedances.
In the first case, every exceedance of $\phi(\beta)$ contributes an exceedance of $\phi(\eta)$; furthermore, by Lemma 4.7, the values $(c,n)$ contribute an additional $t$ exceedances where $t\in\{0,\ldots,s-1\}$ with $t\equiv s-c_{n}\pmod{s}$. Hence each of the $s$ possible values of $t$ occurs exactly once, and the number of possible $\eta$ is given by
(3) $\sum_{t=0}^{s-1}D_{s}(n-1,r-t).$
We now turn to the second case. Let $x$ be the size of $\operatorname{Exc}(\phi(\eta))\cap\{(c,i_{j})\}$, let $y$ be the size of $\operatorname{Exc}(\phi(\eta))\cap\{(c,n)\}$, and let $z$ be the size of $\operatorname{Exc}(\phi(\beta))\cap\{(c,i_{j})\}$. That is, when we change from $\beta$ to $\eta$, we gain $x+y$ exceedances, and we lose $z$, for a net change of $x+y-z$.
By Lemma 4.7,
$x+y-z\equiv(s-c_{n})+(s-c^{\prime}_{j})-(s-c_{j})\equiv 0\pmod{s},$
by our choice of $c^{\prime}_{j}$. Also, since $1\leq x\leq s$, $10\leq y\leq s-$ 1, and $0\leq z\leq s$, we must have $-s+1\leq x+y-z\leq 2s-1$. We conclude that either $x+y-z=0$ or $x+y-z=s$; that is, $\operatorname{exc}(\phi(\beta))\in\{r,r-s\}$.
Let us write $(c,i_{j})=\phi(\eta)^{-1}(s-1,n)$. We know from Lemma 4.7 that $\operatorname{Exc}(\phi(\eta))\cap\{(t,i_{j})\}=\{(t,i_{j})\mid 0\leq t\leq c\}$ and $x=c+1$. The key question is whether $(c,i_{j})$ is an exceedance of $\phi(\beta)$.
If $(c,i_{j})\in\operatorname{Exc}(\phi(\beta))$, then, by Lemma 4.7, we know that $(t,i_{j})\in\operatorname{Exc}(\phi(\beta))$ for $t\leq c$, so $z\geq x$, implying $x+y-z<s$, and hence $x+y-z=0$ and $\operatorname{exc}(\phi(\beta))=r$. For each $\beta$ with $\operatorname{exc}(\phi(\beta))=r$, there are $r$ choices of $(c,i_{j})$.
If $(c,i_{j})\not\in\operatorname{Exc}(\phi(\beta))$, then, again by Lemma 4.7, we know that $(t,i_{j})\not\in\operatorname{Exc}(\phi(\beta))$ for $t\geq c$, so $z<x$, implying $x+y-z>0$, and hence $x+y-z=s$ and $\operatorname{exc}(\phi(\beta))=r-s$. For each $\beta$ with $\operatorname{exc}(\phi(\beta)=r-s$, there are $(n-1)s-(r-s)=sn-r$ choices of $(c,i_{j})$.
Adding these up, the number of ways to obtain $\eta$ where $\phi(\eta)=r$, in our second case where $n$ swaps with some $i_{j}$ for $j<n$, is
(4) $rD_{s}(n-1,r)+(sn-r)D_{s}(n-1,r-s).$
Adding (3) and (4) gives the desired recurrence. ∎
We note that the recurrence (2) has a natural symmetry:
Corollary 4.8.
We have $D_{s}(n,sn-r-1)=D_{s}(n,r)$.
4.1. Relating to Eulerian numbers
We can relate these counts to the Eulerian numbers by adding up sums of $s$ consecutive numbers $D_{s}(n,k)$. For $s=2$, we can also obtain the Eulerian numbers of type B (see Proposition 4.14).
Definition 4.9.
For each $s,n,k$, we define $F_{s}(n,m)$ to be the sum of $s$ consecutive values of $D_{s}(n,m)$:
$F_{s}(n,m)=\sum_{\ell=0}^{s-1}D_{s}(n,m+\ell).$
Proposition 4.10.
The numbers $F_{s}(n,m)$ satsify the recurrence
$F_{s}(n,m)=(m+s)F_{s}(n-1,m)+(sn-m)F_{s}(n-1,m-s).$
with initial conditions $F_{s}(0,m)=1$ when $-(s-1)\leq m\leq 0$ and $F_{s}(0,m)=0$ otherwise.
Proof.
We induct on $n$. For $n=0$, the sum is $1$ if $D_{s}(0,0)$ occurs in the sum and 00 otherwise, as desired.
For $n>0$, we use Theorem 4.3:
(5) $F_{s}(n,m)=\sum_{\ell=0}^{s-1}D_{s}(n,m+\ell)=\sum_{\ell=0}^{s-1}(m+\ell+1)D_{s}(n-1,m+\ell)\\ +\sum_{\ell=0}^{s-1}(sn-m-\ell)D_{s}(n-1,m+\ell-s)+\sum_{\ell=0}^{s-1}\sum_{t=1}^{s-1}D_{s}(n-1,m+\ell-t).$
We would like to rewrite (5) in the form
$F_{s}(n,m)=\sum\alpha_{i}D_{s}(n-1,i).$
What are the coefficients $\alpha_{i}$?
When $i=m+j$, for $10\leq j\leq s-$ 1, then $D_{s}(n-1,i)$ occurs once with coefficient $m+j+1$ (when $\ell=j$) and $s-1-j$ times with coefficient $1$ (when $j+1\leq\ell\leq s-1$). In this case, $\alpha_{i}=m+s$.
When $i=m-s+j$, for $10\leq j\leq s-$ 1, then $D_{s}(n-1,i)$ occurs once with coefficient $sn-m-j$ (when $\ell=j$) and $j$ times with coefficient $1$ (when $10\leq\ell\leq j-$ 1). In this case, $\alpha_{i}=sn-m$.
When $i<m-s$, or $i\geq m+s$, then $\alpha_{i}=0$, Hence, we get
$F_{s}(n,m)$ $=(m+s)\sum_{j=0}^{s-1}D_{s}(n-1,m+j)+(sn-m)\sum_{j=0}^{s-1}D_{s}(n-1,m-s+j)$
$=(m+s)F_{s}(n-1,m)+(sn-m)F_{s}(n-1,m-s).$
If we group these terms in a particular way, we obtain multiples of the Eulerian numbers:
Proposition 4.11.
For all $n$ and $k$,
$F_{s}(n,sk)=s^{n}A(n,k).$
Proof.
For $n=0$, $F_{s}(n,sk)=1$ when $k=0$ and $F_{s}(n,sk)=0$ otherwise.
For $n>0$, by Proposition 4.10,
$F_{s}(n,sk)=s\big{(}(k+1)F_{s}(n-1,sk)+(n-k)F_{s}(n-1,sk-s)\big{)}.$
But, by Proposition 3.3, this same recurrence is satisfied by $s^{n}A(n,k)$. ∎
Corollary 4.12.
The probability that CircularShift succeeds on $s$-suited permutation Wordle in at most $sr$ rounds is the same as the probability that CircularShift succeeds on permutation Wordle in at most $r$ rounds.
When $s=2$, our “2-suited permutations” are more commonly called signed permutations. If we group the terms $D_{2}(n,m)$ differently, we obtain the Eulerian numbers of type B, which count the number of signed permutations with a specified number of type B descents [CG07].
Definition 4.13.
The Eulerian numbers of type B, denoted $B(n,k)$, satisfy the recurrence
(6) $B(n,k)=(2k-1)B(n-1,k)+(2n-2k+1)B(n-1,k-1)$
with initial conditions $B(1,1)=1$ and $B(1,k)=0$ otherwise.
Proposition 4.14.
We have $F_{2}(n,2k-1)=B(n+1,k+1)$.
Proof.
We have $F_{2}(0,2k-1)=1$ when $k=0$ and 00 otherwise, as desired. Also, the recurrence of Proposition 4.10 yields
$F_{2}(n,2k-1)=(2k+1)F_{2}(n-1,2k-1)+(2n-2k+1)F_{2}(n-1,2k-3).$
By (6),
$B(n+1,k+1)=(2k+1)B(n,k+1)+(2n-2k+1)B(n,k).$
These are the same recurrence. ∎
4.2. Other relations
There is another way to describe the counts $D_{s}(n,k)$ in terms of Eulerian numbers.
Definition 4.15.
For $s>0,n\geq 0,j$, let $C_{s}(n,j)$ be the number of ways to place $j$ indistinguishable pebbles in $n$ pots such that no pot has more than $s-1$ pebbles.
The binomial coefficient ($n\choose j$ is $C_{2}(n,j).$ We give a more general recursive formulation for $C_{s}(n,j).$
Proposition 4.16.
For all $s>0$, we have $C_{s}(0,0)=1$, $C_{s}(0,j)=0$ for $j\neq 0$, and
$C_{s}(n,j)=\sum_{t=0}^{s-1}C_{s}(n-1,j-t).$
Proof.
For $n=0$, the only placement available is the empty placement. For $n>0$, we count first the placement of $t$ pebbles in the last pot and the placement of $j-t$ pebbles in the other $n-1$ pots. ∎
Proposition 4.17.
We have the relation
$D_{s}(n,k)=\sum_{j=0}^{k}C_{s}(n,j)A(n,k-j).$
Proof.
Suppose we have a suited $n$-permutation $\pi$, where $\phi(\pi)$ has $k$ exceedances. For each $\ell$, where $\pi(\ell)=(i_{\ell},c_{\ell})$, let $e_{\ell}$ be the number of exceedances arising from the values $\ell$, $\ell+n$, $\ldots$, $\ell+(s-1)n$ in the $sn$-permutation $\phi(\pi)$.
By Lemma 4.7, when $\ell\in\operatorname{Exc}(\pi)$, then $1\leq e_{\ell}\leq s$. When $\ell\not\in\operatorname{Exc}(\pi)$, then $10\leq e_{\ell}\leq s-$ 1. We can thus write $e_{\ell}=a_{\ell}+b_{\ell}$, where $a_{\ell}$ is $1$ when $\ell$ is an exceedance and 00 when it is not, and $b_{\ell}$ takes all possible values $10\leq b_{\ell}\leq s-$ 1.
If $\pi$ has $k-j$ exceedances, then $\sum a_{\ell}=k-j$, and we need $\sum b_{\ell}=j$. The result follows. ∎
5. Open questions
We have described the game permutation Wordle, and the generalization suited permutation Wordle. We have also described the strategy CircularShift, which we believe solves both games optimally.
The natural open question is whether CircularShift really is optimal for permutation Wordle, or for suited permutation Wordle. Does it dominate all other strategies, in the sense that, for all $r$, it wins more often within $r$ rounds than any other strategy? If not, is it optimal in some more limited sense (e.g., average number of rounds)? Suppose we allow the guesser to guess arbitary strings of numbers, as in Li–Zhu [LZ24]: does that provide a more efficient strategy?
One can also ask about other generalizations. The surprising thing about permutation Wordle is the appearance of $A(n,k)$ and $B(n,k)$ in the analysis; are there other games where other generalizations of the Eulerian numbers naturally arise?
The authors are indebted to Alex Miller for making the connection with exceedances and for proposing the signed variant. We would also like to thank Tim Chow, Keith Frankston, Ben Howard, Danny Scheinerman, and Jeff VanderKam for helpful discussions.