### v2

# Permutation wordle

Abstract.
We introduce a guessing game, permutation Wordle, in which a guesser attempts to recover a hidden permutation in $S_{n}$. In each round, the guesser guesses a permutation (using information from previous rounds) and is told which entries of that permutation are correct. We describe a natural guessing strategy, which we believe to be optimal. We show that the number of permutations this strategy solves in $k+1$ rounds is the Eulerian number $A(n,k)$.
We also describe an extension to $s$-colored permutations: the setter chooses a permutation in $S_{n}$ and also a coloring of $[n]$ using $s$ colors. We generalize our strategy, give a recurrence for the number of $s$-colored permutations solved in $k+1$ rounds, and relate these numbers both to the Eulerian numbers and to higher-order Eulerian numbers.
1. Introduction
Consider the following game: A setter chooses a permutation on $n$ objects, say, cards labeled $1$ through $n$. Your job is to guess this hidden permutation. To make a guess, you arrange the $n$ cards, and the setter tells you which ones are in the correct place; you win if all $n$ are correct. What is the optimal strategy? How many guesses do you expect to need?
We call this game “permutation Wordle”, because the guesser’s feedback is similar to that in Josh Wardle’s popular online game Wordle. (For connections to other games, and to prior analysis of those games, see Section 2.) We propose a simple strategy, which we call CircularShift: the initial guess is an arbitrary permutation (e.g., the identity), and each successive guess cyclically shifts any incorrectly placed cards one step to the right (skipping over and leaving in place the correctly placed cards). We prove the following:
Theorem 1.1.
The number of permutations solved in $k+1$ guesses by CircularShift is the Eulerian number $A(n,k)$: the number of permutations in $S_{n}$ with $k$ excedances.
The Eulerian numbers come up in a variety of contexts in combinatorics; we discuss several definitions, and prove Theorem 1.1, in Section 3.
We conjecture that CircularShift is optimal in the following sense: Suppose the setter chooses a permutation uniformly from $S_{n}$. Given a strategy ${\mathcal{S}}$, for each $r$, let $P_{r}({\mathcal{S}})$ denote the probability that ${\mathcal{S}}$ guesses the correct permutation within $r$ guesses.
Conjecture 1.2.
$P_{r}(\textnormal{{CircularShift}})$ attains the maximal value of $P_{r}$ for all $r\geq 1$.
We also consider a game we call “multicolor permutation Wordle”; imagine that there is a deck of $sn$ cards labeled $1$ through $n$ in each of $s$ different colors, or more generally a deck of $\sum s_{i}$ cards where for each $i$ there are $s_{i}$ differently colored cards labeled by $i$. The game is harder because some of your guessed cards may not be in the setter’s arrangement at all; you may have the wrong color.
In Section 4 we describe and analyze a generalization of the CircularShift strategy, and we show that Eulerian numbers continue to play a role. For example, when $s=2$, the number of $2$-colored permutations solved in $12k+$ 1 or $22k+$ 2 rounds is $2^{n}A(n,k)$. However, adding the number solved in $2k$ or $12k+$ 1 rounds gives the Eulerian numbers of type B. Thus, permutation Wordle provides a surprising connection between these two sequences. We derive recurrences and generating functions for the general problem of $(s_{1},\ldots,s_{n})$-colored permutations.
In Section 5 we focus on $s$-colored permutations. We show that, for larger $s$, taking consecutive sums of $s$ rounds at a time recovers higher-order Eulerian numbers and some further generalizations. We conclude in Section 6 with some open questions.
2. Guessing games
In the classic game Mastermind, invented by Mordecai Meirowitz, the setter chooses a “code”: a string of length $n$ from an alphabet of size $m$ (which we identify as $[m]=\{1,\ldots,m\}$). In each round, the guesser tries a string in $[m]^{n}$, and the setter responds with two numbers: the number of entries that are correct (right number, right place) and the number of additional entries that are present elsewhere in the code (right number, wrong place). Donald Knuth [Knu77] showed that, for the original game ($n=4$ and $m=6$), five rounds are sufficient for the guesser to guess the code. Many others have evaluated the game for various values of $n$ and $m$. The typical emphasis is on a worst-case analysis; one can imagine the setter adversarially altering the code to be consistent with all information presented thus far, with the goal of maximizing the number of rounds.
One Mastermind variant is “permutation Mastermind”, in which $m=n$ and both the setter’s code and the guesser’s guesses must contain each value exactly once. (In this variant, the setter responds with the number $c$ of correct entries, since the number in the wrong place is simply $n-c$.) El Ouali et al. [EOGSS18] have shown that $n\log_{2}n+O(n)$ queries suffice.
In 2020, a new online game became popular: Wordle, by Josh Wardle. The setup is the same as Mastermind, but the feedback is different: the setter indicates which entries are correct and in the correct place, and which additional entries are present in the code elswhere. [^1] The rules for handling cases where a value occurs multiple times in the code or the guess are irrelevant to our permutation problem. In the original Wordle, the alphabet is the letters A–Z, and the code and guesses are required to be 5-letter words in English. There are numerous other variants, including a recent one called “OrderUp” [Wag24] in which the code and guesses are 7-long permutations.
Li and Zhu [LZ24] have analyzed a generalized version of Wordle, with a code from the alphabet $[m]^{n}$. Among other things, they show that, if $m=n$ and the code is required to use each value exactly once, then the guesser needs $n$ guesses in the worst case; this is true even if the guesses are are allowed to have values multiple times. They give a simple strategy with a matching upper bound of $n$ guesses for the general case $m=n$: the first $n-1$ guesses are the strings $1^{n}$, $2^{n}$, $\ldots$, $(n-1)^{n}$, at which point the guesser knows the code. We show that, in fact, the guesser can guess an $n$-long permutation in at most $n$ guesses even when the guesses are required to be permutations.
As with the Mastermind analyses, Li and Zhu emphasize worst-case behavior; any average-case analysis would need to consider a probability distribution over codes. As we will see below, in the case of permutations, all codes are of equal difficulty; we may as well assume the setter chooses a random code. Therefore, we consider average-case behavior. More generally, for each $r\leq n$, we consider the probability that the guesser succeeds in $r$ guesses.
3. Eulerian numbers and CircularShift
We describe a simple strategy, CircularShift, which solves permutation Wordle in at most $n$ guesses. As noted above, Li and Zhu [LZ24] have previously shown that $n$ guesses are necessary. In our setting, we investigate average-case behavior.
Recall that the setter chooses a permutation $\sigma\in S_{n}$; let $\{\gamma_{r}\}$ for $r\geq 1$ denote the guesser’s permutations. The feedback in round $r$ is the set $\{i\mid\gamma_{r}(i)=\sigma(i)\}$, or, equivalently, the set of fixed points of $\sigma\gamma_{r}^{-1}$. If the setter had instead chosen $\sigma\tau$, and the guesser’s $r$th guess were $\gamma_{r}\tau$, the feedback sequence would be the same. The guesser could choose a uniform random $\tau$, and multiply each guess by $\tau$ throughout. Therefore, we may assume that the setter chooses $\sigma$ uniformly at random, and, for simplicity, we let $\gamma_{1}$ be the identity permutation.
We discuss average-case behavior: we compute the probability of solving permutation Wordle in exactly $r$ guesses for various values of $r$. For example, consider the case $n=3$. Assume the setter’s $\sigma$ is chosen randomly, and that our first guess $\gamma_{1}$ is the identity permutation $(1,2,3)$; we then learn the set of fixed points of $\sigma$. If $\sigma$ is the identity, we are immediately done; if $\sigma$ has one fixed point, we can get $\sigma$ in round 2. If $\sigma$ is one of the two 3-cycles, then our best strategy is to guess a 3-cycle in round 2; we will either be correct, or will need a third guess. Overall, the probability of finding $\sigma$ in one round is $1/6$; in two rounds, $4/6$; in three rounds, $1/6$.
The ananlysis for $n=4$ is more complicated. Again, let $\gamma_{1}$ be the identity. If $\sigma$ is the identity, we finish in one round; if it has two fixed points, we need two rounds; if it has one fixed point, we need two or three rounds. When $\sigma$ is one of the nine derangements, the optimal strategy turns out to be to let $\gamma_{2}$ be a $4$-cycle. If $\sigma=\gamma_{2}$, we are done in two rounds; if $\sigma\gamma_{2}^{-1}$ has any fixed points, we can deduce $\sigma$ in the third round; otherwise, $\sigma\in\{\gamma_{2}^{2},\gamma_{2}^{3}\}$, and we need three or four rounds.
Overall, for $n=4$, the probability of finishing in one round is $1/24$; in two rounds, $11/24$; in three rounds, $11/24$; in four rounds, $1/24$. Equivalently, the probabilities $P_{r}$ of finishing it at most $r$ rounds are $1/24$, $1/2$, $23/24$, $1$. Experiments confirm that for $n=4$ this strategy dominates all others in the following sense: if we compute the probabilities $P^{\prime}_{r}$ for any other strategy, then $P_{r}\geq P^{\prime}_{r}$ for all $r$.
3.1. The strategy CircularShift
We now extend this strategy to larger $n$. We describe a deterministic approach to generating the guesses $\gamma_{r}$; as noted above, for a randomized version, we could choose a permutation $\tau$ and guess $\gamma^{\prime}_{r}=\gamma_{r}\tau$.
Definition 3.1 (CircularShift).
Let $\gamma_{1}=(1,2,\ldots,n)$. Suppose that, in round $r$, we submit guess $\gamma_{r}$ and receive feedback $U_{r}$, the set of values on which $\gamma_{r}$ agrees with the setter’s permutation $\sigma$. If $U_{r}=[n]$, we are done in $r$ rounds. Otherwise, we define $\gamma_{r+1}$ as follows: write $[n]\setminus U_{r}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$, and assign
$\gamma_{r+1}(i):=\begin{cases}\gamma_{r}(i)&\text{if $i\in U_{r}$,}\\ \gamma_{r}(a_{j-1})&\text{if $i=a_{j}$ for some $j>1$,}\\ \gamma_{r}(a_{\ell})&\text{if $i=a_{1}$.}\end{cases}$
Let $D(n,r)$ denote the number of permutations for which CircularShift succeeds in exactly $r$ rounds.
1 2 3 4 5 6 7 8 9
$\gamma_{1}$ 1 2 3 4 5 6 7 8 9
$\gamma_{2}$ 8 2 1 3 5 4 6 7 9
$\gamma_{3}$ 7 2 8 1 5 3 4 6 9
$\gamma_{4}$ 7 2 4 8 5 3 1 6 9
Table 1. The deterministic strategy permutation Wordle guessing the permutation $\sigma=(7,2,4,8,5,3,1,6,9)$. Correct values are shown in boldface; the underlines will be discussed in Section 3.3. The strategy succeeds in four rounds.
We work out an example in Table 1. In the first round, we guess the identity permutation and get three values correct. In round 2, we cyclically shift the remaining six values; none of them are correct, so we shift them again in round 3. At this point, we have six correct, and in round 4 we fix the remaining values.
The probability of succeeding in at most $r$ rounds is $\sum_{k=1}^{r}D(n,k)/n!$. The claim of Conjecture 1.2 is that these probabilities are all at least as large for CircularShift as for any other strategy. We have proven this for $n\leq 5$.
3.2. Eulerian numbers
The Eulerian numbers occur frequently in the combinatorics of permutations. One common definition involves counting the number of descents in a permutation: that is, the number of values $i$ for which $\pi(i)>\pi(i+1)$. For our purposes, it will be more convenient to use a formulation in terms of excedances.
Definition 3.2.
For $\pi\in S_{n}$, an excedance is a value $i$ for which $\pi(i)>i$. We let $\operatorname{Exc}(\pi)$ denote the set of excedances: $\operatorname{Exc}(\pi):=\{i\mid\pi(i)>i\}$. We write $\operatorname{exc}(\pi)=\#\operatorname{Exc}(\pi)$.
For $0\leq k<n$, the Eulerian number $A(n,k)$ is #$\#\{\pi\in S_{n}\mid\operatorname{exc}(\pi)=k\}$.
For example, when $n=3$, $A(3,0)=1$ (the only such permutation is $(1,2,3)$) and $A(3,2)=1$ (the only such permutation is $(2,3,1)$); we then have $A(3,1)=4$.
Proposition 3.3.
The counts $A(n,k)$ satisfy the recurrence
(1) $A(n,k)=(k+1)A(n-1,k)+(n-k)A(n-1,k-1)$
with initial conditions $A(0,0)=1$ and $A(0,k)=0$ for $k\neq 0$.
Proof.
We induct on $n$. The unique element of $S_{0}$ has 00 excedances, verifying the initial condition.
Every permutation $\pi\in S_{n}$ can be uniquely decomposed in the form $\alpha\beta$, where $\beta$ fixes $n$ and $\alpha$ is either the identity, or swaps $n$ with some value $j$. We view $\beta$ as an element of $S_{n-1}$.
How can $\pi$ have exactly $k$ excedances? There are two possibilities. First, $\beta$ could have $k$ excedances, and $\alpha$ is either the identity, or swaps $n$ with one of the $k$ places in $\operatorname{Exc}(\beta)$. Second, $\beta$ could have $k-1$ excedances, and $\alpha$ swaps $n$ with one of the $n-k$ places not in $\operatorname{Exc}(\beta)$. This corresponds to the desired recurrence. ∎
It is straightforward to show that, if we had defined $A(n,k)$ in terms of the number of permutations with $k$ descents, the same recurrence would apply. Alternatively, Foata and Schützenberger [FS70] givee an explicit bijection between permutations with $k$ excedances and those with $k$ descents.
Note that the recurrence implies the reflective symmetry $A(n,k)=A(n,n-k-1)$.
The Eulerian numbers show up in a variety of other ways; for example, if $Y_{i}$ are uniform random variables in the interval $(0,1)$, and we define $X=\left\lfloor{\sum_{i=1}^{n}Y_{i}}\right\rfloor$, then the probability that $X=k$ is given by $A(n,k)/n!$. We now show that these numbers also arise in permutation Wordle.
3.3. The main theorem
We connect the performance of CircularShift with excedances.
Theorem 3.4.
The number of guesses it takes CircularShift to guess a permutation $\sigma$ is $1+\operatorname{exc}(\sigma)$.
Note that Theorem 1.1 follows immediately from Theorem 3.4 and from Definition 3.2.
Proof of Theorem 3.4.
Let $\sigma$ be the setter’s permutation; we run CircularShift, producing a series of guesses $\gamma_{1}$, …, $\gamma_{r}$ for some $r$ with $\gamma_{r}=\sigma$. Define the set $R_{k}$ to be the values that $\gamma_{k}$ places further right than $\sigma$ does:
$R_{k}=\{i\mid\gamma_{k}^{-1}(i)>\sigma^{-1}(i)\}.$
(The numbers in $R_{k}$ are underlined in Table 1.) We again let $U_{k}$ denote the values that are already fixed; that is, those $i$ with $\gamma_{k}^{-1}(i)=\sigma^{-1}(i)$.
How does $R_{k}$ change as $k$ increases? Let $\ell$ denote the rightmost entry in $\gamma_{k}$ not in $U_{k}$; then $\ell$ is to the right of its position in $\sigma$, so $\ell\in R_{k}$. In $\gamma_{k+1}$, this value $\ell$ will cycle around to the first available position (that is, anything to its left is in $U_{k}$), so $\ell\not\in R_{k+1}$.
Any other values in $R_{k}$ will move further right, and will remain in $R_{k+1}$. Any values in $U_{k}$ will remain where they were. Finally, any other values (those $i$ with $\gamma_{k}^{-1}(i)>\sigma^{-1}(i)$) will move right one step, but cannot move past their location in $\sigma$, so they will not be in $R_{k+1}$. We conclude that $R_{k+1}=R_{k}\setminus\{\ell\}$, so #$\#R_{k+1}=\#R_{k}-1$.
CircularShift terminates when $R_{r}$ is empty (that is, when $\gamma_{r}=\sigma$), so the number of rounds is $1+\#R_{1}$. Finally, $i\in R_{1}$ if and only if $i>\sigma^{-1}(i)$, or $\sigma^{-1}(i)\in\operatorname{Exc}(\sigma)$. ∎
Corollary 3.5.
The expected number of rounds for CircularShift is $(n+1)/2$.
Proof.
This follows immediately from the symmetry of the Eulerian numbers: $A(n,k)=A(n,n-k-1)$. ∎
This leads us to a weaker version of Conjecture 1.2:
Conjecture 3.6.
For any strategy for permutation Wordle, the average number of rounds is at least $(n+1)/2$.
4. Multicolor permutation Wordle
We now turn to the “multicolor” version of permutation Wordle. A multicolor permutation is a permutation in $S_{n}$ together with a coloring $c\colon[n]\to\{0,1,\ldots\}$. An $(s_{1},\ldots,s_{n})$-colored permutation is one where $0\leq c(i)<s_{i}$ for all $i$, and an $s$-colored permutation is one where $0\leq c(i)<s$. We focus on these cases where $c(i)$ is bounded so that we can talk about a uniform random $(s_{1},\ldots,s_{n})$-colored or $s$-colored permutation.
The setter’s $\sigma$ with coloring $\kappa$, and the guesses $\gamma_{k}$ with colorings $c_{k}$, are each now a multicolor permutation. The feedback [^2] A direct application of the Wordle rules would be that a color/number combo is labeled green if it is the right place, and yellow if it occurs in the permutation but in a different place, but this leads to a less interesting math problem. will be as before: we are told the list of $j$ where our guess agrees with $\sigma$.
As with “monochrome” permutation Wordle, we can assume that the setter chooses $\sigma$ and $\kappa$ uniformly at random, and that our first guess $(\gamma_{1},c_{1})$ is the identity permutation $\gamma_{1}$ and $c_{1}(i)=0$ for all $i$.
4.1. CircularShift for multicolor permutation Wordle
Definition 4.1 (CircularShift for multicolor permutation Wordle).
Let $\gamma_{1}=(1,2,\ldots,n)$, and let $c_{1}\equiv 0$. Suppose that, in round $r$, we submit guess $(\gamma_{r},c_{r})$ and receive feedback $U_{r}$, the set of values $k$ on which $\gamma_{r}(k)=\sigma(k)$ and $c_{r}(\sigma(k))=\kappa(\sigma(k))$. If $U_{r}=[n]$, we are done in $r$ rounds. Otherwise, we define $\gamma_{r+1}$ and $c_{r+1}$ as follows: write $[n]\setminus U_{r}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$. Let
$\gamma_{r+1}(i):=\begin{cases}\gamma_{r}(i)&\text{if $i\in U_{r}$,}\\ \gamma_{r}(a_{j-1})&\text{if $i=a_{j}$ for some $j>1$,}\\ \gamma_{r}(a_{\ell})&\text{if $i=a_{1}$.}\end{cases}$
Also,
$c_{r+1}(i):=\begin{cases}c_{r}(i)+1&\text{$i=\gamma_{r+1}(a_{j})$ for some $j>1$ and $\gamma_{r+1}(a_{j-1})<a_{j}\leq\gamma_{r+1}(a_{j})$,}\\ c_{r}(i)+1&\text{$i=\gamma_{r+1}(a_{1})$ and either $a_{1}\leq\gamma_{r+1}(a_{1})$ or $\gamma_{r+1}(a_{\ell})<a_{1}$,}\\ c_{r}(i)&\text{otherwise}.\end{cases}$
In the monochrome case, CircularShift is the same as in Definition 3.1; we rotate the values of $\gamma_{r}$ just as before, and we start with all colors equal to 00. When a value $i$ with color $a$ has looped all the way around to its starting position, we know the color must have been incorrect, so we increment it to $a+1$.
We next prove a multicolor analogue of Theorem 3.4.
Theorem 4.2.
The number of guesses it takes CircularShift to guess a permutation $\sigma$ with coloring $\kappa$ is given by
$1+\operatorname{exc}(\sigma)+\sum_{i=1}^{n}\kappa(i).$
The idea of the proof is to introduce quantities $Z_{r,i}$ representing the number of times after round $r$ that the value $i$ wraps around before we finish solving. In each round, one value ($i=\gamma_{r}(a_{\ell})$) wraps around, so the total $\sum_{i}Z_{r,i}$ decreases by $1$. We are done when it reaches zero.
Proof.
First, we show CircularShift eventually succeeds on any permutation $\sigma$ with coloring $\kappa$. For each $i$, if we have not yet gotten $i$ correct, then $i$ will continue to cycle around. Each time $i$ passes its starting location, we increase the value of $c_{r}(i)$, until $c_{r}(i)=\kappa(i)$. Since $\sigma(i)$ is still a location we have not filled, $i$ cannot skip past it, so we eventually try $i$ in the right location with the right color.
As in the proof of Theorem 3.4, we need a quantity which decreases by $1$ in each round and which is 00 when we have the right answer. Suppose that, in round $r$, we have some guess $\gamma_{r}$ with coloring $c_{r}$. For each $i$, let $\delta(i)=1$ when $i$ occurs further to left in $\sigma$ than in the identity; that is, $i>\sigma^{-1}(i)$, or $\sigma^{-1}(i)$ is an excedance of $\sigma$. Similarly, let $\epsilon_{r}(i)=1$ when $i>\gamma_{r}^{-1}(i)$. Now, for each $i$, let
$Z_{r,i}=\kappa(i)-c_{r}(i)+\delta(i)-\epsilon_{r}(i)$
and let $Z_{r}=\sum_{i=1}^{n}Z_{r,i}$. It is clear that $Z_{1}$ is equal to $\operatorname{exc}(\sigma)+\sum_{i}\kappa(i)$. Also, if $\gamma_{r}=\sigma$ and $c_{r}=\kappa$, we must have $Z_{r}=0$.
How can we have $Z_{r+1,i}\neq Z_{r,i}$? Either $c_{r+1}(i)\neq c_{r}(i)$, or $\epsilon_{r+1}(i)\neq\epsilon_{r}(i)$. There are only two ways this can happen: first, we could have $i=\gamma_{r}(a_{j})$ for some $j<\ell$, and $a_{j}<i\leq a_{j+1}$. Second, we could have $i=\gamma_{r}(a_{\ell})$; this second case will happen for exactly one $i$ each time.
In the first case, $\gamma_{r+1}^{-1}(i)>i>\gamma_{r}^{-1}(i)$, so $\epsilon_{r}(i)=0$ and $\epsilon_{r+1}(i)=1$. In this situation $c_{r+1}(i)=c_{r}(i)+1$. Thus $Z_{r+1,i}=Z_{r,i}$.
In the second case, there are three possibilities: if $i>a_{\ell}\geq a_{1}$, then $\epsilon_{r+1}(i)=\epsilon_{r}(i)=1$ and $c_{r+1}(i)=c_{r}(i)+1$. If $a_{\ell}\geq a_{1}\geq i$, then $\epsilon_{r+1}(i)=\epsilon_{r}(i)=0$ and $c_{r+1}(i)=c_{r}(i)+1$. If $a_{\ell}\geq i>a_{1}$, then $\epsilon_{r+1}(i)=0$ and $\epsilon_{r}(i)=1$, and $c_{r+1}(i)=c_{r}(i)$. In each case, $Z_{r+1,i}=Z_{r,i}-1$. Since this is the only $i$ for which $Z_{r+1,i}\neq Z_{r,i}$, we have $Z_{r+1}=Z_{r}-1$.
We showed above that, for some $r$, $\gamma_{r}=\sigma$ and $c_{r}=\kappa$, at which point $Z_{r}=0$. This happens when $r-1=Z_{1}=\operatorname{exc}(\sigma)+\sum_{i}\kappa(i)$. ∎
Conjecture 4.3.
For each $r$, the probability that CircularShift solves a uniformly chosen $(s_{1},\ldots,s_{n})$-colored permutation within $r$ rounds is optimal, that is, it is at least as large as for any other strategy.
4.2. Multicolor recurrences and generating functions
We want to understand the expected behavior of CircularShift on multicolor permutations with various numbers of colors.
Definition 4.4.
We let ${D_{(s_{1},\ldots,s_{n})}}(n,k)$ denote the number of $(s_{1},\ldots,s_{n})$-colored permutations of length $n$ for which CircularShift succeeds in exactly $k+1$ rounds. We let $D_{s}(n,r)$ denote the number of $s$-colored permutations for which CircularShift succeeds in exactly $r+1$ rounds.
There is a well known generating function representation of the Eulerian numbers: [^3] Theorem 4.6, setting all $s_{i}=1$, yields a proof that the coefficients given by (2) satisfy the recurrence (1).
(2) $\frac{\sum_{k}A(n,k)t^{k}}{(1-t)^{n+1}}=\sum_{j}(j+1)^{n}t^{j}.$
Theorem 4.5.
(3) $\frac{\sum_{k}{D_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}}{(1-u)\prod_{\ell}(1-u^{s_{\ell}})}=\sum_{j}(j+1)^{n}u^{j}.$
Proof.
By Theorem 4.2,
$\sum_{k}{D_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}$ $=\sum_{\sigma,\kappa}u^{\operatorname{exc}(\sigma)+\kappa(1)+\cdots+\kappa(n)}$
$=\left(\sum_{\sigma}u^{\operatorname{exc}(\sigma)}\right)\left(\sum_{\kappa}\prod_{\ell=1}^{n}u^{\kappa(\ell)}\right)$
$=\left(\sum_{k}A(n,k)u^{k}\right)\prod_{\ell=1}^{n}(1+u+\cdots+u^{s_{\ell}-1}),$
since $A(n,k)$ counts excedances. Using (2),
$\sum_{k}{D_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}$ $=\left(\sum_{j}(j+1)^{n}u^{j}\right)(1-u)^{n+1}\prod_{\ell=1}^{n}\frac{1-u^{s_{\ell}}}{1-u}.$
The result follows. ∎
From the generating function (3) we can derive a recurrence for ${D_{(s_{1},\ldots,s_{n})}}(n,k)$. Note that ${(s_{1},\ldots,\widehat{s_{\ell}},\ldots,s_{n})}$ denotes the $(n-1)$-long sequence obtained from $(s_{1},\ldots,s_{n})$ by removing $s_{\ell}$.
Theorem 4.6.
The counts ${D_{(s_{1},\ldots,s_{n})}}(n,k)$ satisfy the recurrence
[…]
with initial conditions $D_{()}(0,0)=1$ and $D_{()}(0,k)=0$ for $k\neq 0$.
The proof uses standard generating function techniques; see Appendix A.
Theorems 4.5 and 4.6 are both simpler in the $s$-colored case:
Theorem 4.7.
$\frac{\sum_{k}D_{s}(n,k)u^{k}}{(1-u)(1-u^{s})^{n}}=\sum_{j}(j+1)^{n}u^{j}.$
Theorem 4.8.
The counts $D_{s}(n,k)$ satisfy the recurrence
(5) $D_{s}(n,k)=(k+1)D_{s}(n-1,k)+(sn-k)D_{s}(n-1,k-s)+\sum_{\ell=1}^{s-1}D_{s}(n-1,k-\ell)$
with initial conditions $D_{s}(0,0)=1$ and $D_{s}(0,k)=0$ for $k\neq 0$.
Note that, when $s=1$, this is the recurrence (1) for the Eulerian numbers.
The recurrence (5) has a natural symmetry:
Corollary 4.9.
We have $D_{s}(n,sn-r-1)=D_{s}(n,r)$.
4.3. Sums of consecutive values
We can find patterns in the sums of consecutive values of these sequences $D_{s}(n,r)$. For example, consider adding pairs of numbers from $D_{2}(n,r)$; the case $n=4$ is shown in Figure 1. If we start with the first two numbers, then the next two, and so on, we obtain multiples of the Eulerian numbers: the sequence is $2^{n}A(n,r)$. However, if we instead start with just one number, then two at a time, we obtain the Eulerian numbers of type B [CG07]. We now develop a recurrence for these sums of consecutive values; in Section 5 we show that we recover the higher-order Eulerian numbers [Ste94].
11561115115611511617617616176230761
Figure 1. Two ways to group the sequence $D_{2}(4,k)$.
Definition 4.10.
For each $s,n,k$, we define $F_{s}(n,k)$ to be the sum of $s$ consecutive values of $D_{s}(n,r)$:
$F_{s}(n,k)=\sum_{\ell=0}^{s-1}D_{s}(n,k-\ell).$
Theorem 4.11.
(6) $\frac{\sum_{k}F_{s}(n,k)u^{k}}{(1-u^{s})^{n+1}}=\sum_{j}(j+1)^{n}u^{j}.$
Proof.
From the definition, $\sum_{k}F_{s}(n,k)u^{k}=(1+\cdots+u^{s-1})\sum_{k}D_{s}(n,k)u^{k}$. The result follows immediately from Theorem 4.7. ∎
Theorem 4.12.
The numbers $F_{s}(n,k)$ satsify the recurrence
$F_{s}(n,k)=(k+1)F_{s}(n-1,k)+(s(n+1)-k-1)F_{s}(n-1,k-s).$
with initial conditions $F_{s}(0,m)=1$ when $10\leq m\leq s-$ 1 and $F_{s}(0,m)=0$ otherwise.
Again, a proof can be found in Appendix A.
Remark 4.13.
The sequences ${D_{(s_{1},\ldots,s_{n})}}(n,\cdot)$ and $F_{s}(n,\cdot)$ are log-concave. Note that ${D_{(s_{1},\ldots,s_{n})}}(n,\cdot)$ is the iterated convolution of $A(n,\cdot)$ with sequences $(1,\ldots,1)$ of length $s_{\ell}$ for $1\leq\ell\leq n$, and $F_{s}(n,\cdot)$ is the convolution of $D_{s}(n,\cdot)$ with $(1,\ldots,1)$. The convolution of two log-concave sequences is log-concave [KG71].
5. Higher-order Eulerian numbers
Theorem 4.12 implies that multicolor permutation Wordle recovers multiples of the Eulerian numbers.
Proposition 5.1.
For all $n$ and $k$,
$F_{s}(n,sk+s-1)=s^{n}A(n,k).$
Proof.
For $n=0$, $F_{s}(n,sk+s-1)=1$ when $k=0$ and $F_{s}(n,sk+s-1)=0$ otherwise.
For $n>0$, by Theorem 4.12,
$F_{s}(n,sk+s-1)$ $=(sk+s)F_{s}(n-1,sk+s-1)+(s(n+1)-sk-s)F_{s}(n-1,sk-1)$
$=s\left[(k+1)F_{s}(n-1,sk+s-1)+(n-k)F_{s}(n-1,s(k-1)+s-1)\right].$
By Proposition 3.3, this same recurrence is satisfied by $s^{n}A(n,k)$. ∎
Corollary 5.2.
The probability that CircularShift succeeds on $s$-colored permutation Wordle in at most $sr$ rounds is the same as the probability that CircularShift succeeds on permutation Wordle in at most $r$ rounds.
5.1. $z$-excedances
To describe higher-order Eulerian numbers, we follow (and slightly generalize) the approach of Steingrímsson [Ste94].
Definition 5.3.
For an $s$-colored permutation $\pi$, a $z$-excedance is a value $j$ for which either $\pi(j)>j$, or for which $\pi(j)=j$ and $c_{i}\geq z$. Equivalently, it is a value $j$ where $(i_{j},c_{j})\succ(i,z-1)$ lexicographically.
For $1\leq z\leq s$, we let $E_{s,z}(n,k)$ denote the number of $s$-colored permutations of length $n$ with $k$ $z$-excedances.
Steingrímsson analyzes the case $z=1$ and derives the recurrence
$E_{s,1}(n,k)=(sk+1)E_{s,1}(n-1,k)+(s(n-k)+s-1)E_{s,1}(n-1,k-1)$
and the generating function
(7) $\frac{\sum_{k}E_{s,1}(n,k)t^{k}}{(1-t)^{n+1}}=\sum(sj+1)^{n}t^{j}.$
We generalize these results to larger values of $z$:
Theorem 5.4.
The numbers $E_{s,z}(n,k)$ satisfy the recurrence
$E_{s,z}(n,k)=(sk+z)E_{s,z}(n-1,k)+(s(n-k)+s-z)E_{s,z}(n-1,k-1)$
with initial conditions $E_{s,z}(0,k)=1$ when $k=0$ and $E_{s,z}(0,k)=0$ for $k\neq 0$.
The proof follows the same lines as that of Proposition 3.3.
Proof.
We induct on $n$; the unique element of $S_{0}$ has 00 $z$-excedances, verifying the initial condition.
Every $s$-colored permutation $\pi$ in $S_{n}$ can be uniquely constructed as follows: start with an $s$-colored permutation $\beta\in S_{n-1}$. Append $(n,c)$ for some color $c$, and then either leave it there, or swap it with $(i,c_{i})$ for some $i<n$.
How can $\pi$ have exactly $k$ $z$-excedances? There are two possibilities. First, $\beta$ could have $k$ $z$-excedances, and either we swap $n$ with one of the $k$ $z$-excedances (in which case any color $c$ works), or we leave $n$ in place and $c<z$. Given a permutation $\beta$ with $k$ $z$-excedances, this yields $sk+z$ permutations $\pi$.
Second, $\beta$ could have $k-1$ $z$-excedances, and either we swap $n$ with one of the $n-k$ non-$z$-excedances (in which case any color $c$ works), or we leave $n$ in place and $c\geq z$ (giving one more $z$-excedance). Given a permutation $\beta$ with $k-1$ $z$-excedances, this yields $s(n-k)+s-z$ permutations $\pi$.
The sum of these two cases yields the desired recurrence. ∎
Note that $E_{s,s}(n,k)$ is simply $s^{n}A(n,k)$; since no color can exceed $s-1$, the colors play no role and we simply count excedances. When $z=s$, the recurrence of Theorem 5.4 simplifies to
$E_{s,s}(n,k)=s(k+1)E_{s,s}(n-1,k)+s(n-k)E_{s,s}(n-1,k-1),$
which is the recurrence for $F_{s}(n,sk+s-1)$ discussed in Proposition 5.1. We can apply similar logic to any value of $z$.
Proposition 5.5.
For $1\leq z\leq s$, we have
$F_{s}(n,sk+z-1)=E_{s,z}(n,k).$
Proof.
When $n=0$, $F_{s}(n,sk+z-1)$ is $1$ when $k=0$ and 00 otherwise, as desired.
For $n>0$, by Theorem 4.12,
$F_{s}(n,sk+z-1)=(sk+z)F_{s}(n-1,sk+z-1)+(s(n-k)+s-z)F_{s}(n-1,s(k-1)+z-1).$
By Proposition 5.4, this is the recurrence satisfied by $E_{s,z}(n,k)$. ∎
Proposition 5.6.
$E_{ds,dz}(n,k)=d^{n}E_{s,z}(n,k).$
Proof.
The proof is the same as that of Proposition 5.1, except that we pull out a factor of $d$, rather than $s$, each time we apply the recurrence. ∎
Remark 5.7.
Since $E_{s,z}(n,\cdot)$ is a subsequence of the log-concave $F(n,\cdot)$ (see Remark 4.13), $E_{s,z}(n,\cdot)$ is also log-concave.
Theorem 5.8.
For $1\leq z\leq s$,
$\frac{\sum_{k}E_{s,z}(n,k)t^{k}}{(1-t)^{n+1}}=\sum_{j}(sj+z)^{n}t^{j}.$
Note that, when $z=1$, this recovers Steingrímsson’s (7).
Proof.
By Theorem 4.11,
$\sum_{k}F_{s}(n,k)u^{k}=(1-u^{s})^{n+1}\sum_{j}(j+1)^{n}u^{j}.$
On each side, consider only terms corresponding to $u^{k}$ with $k\equiv z-1\pmod{s}$:
$\sum_{k}F_{s}(n,sk+z-1)u^{sk+z-1}=(1-u^{s})^{n+1}\sum_{j}(sj+z)^{n}u^{sj+z-1}.$
We divide both sides by $u^{z-1}$ and set $t=u^{s}$: with Proposition 5.5, this proves the identity. ∎
We can also prove a version of the classic identity
$\sum_{n=0}^{\infty}\frac{x^{n}}{n!}\left(\sum_{k=0}^{n}A(n,k)t^{k}\right)=\frac{t-1}{t-e^{x(t-1)}}$
and of Steíngrimsson’s similar identity for $E_{s,1}(n,k)$ [Ste94].
Corollary 5.9.
$\sum_{n=0}^{\infty}\frac{x^{n}}{n!}\left(\sum_{k=0}^{n}E_{s,z}(n,k)t^{k}\right)=\frac{(1-t)e^{zx(1-t)}}{1-te^{sx(1-t)}}.$
Using Theorem 4.11, we can do the same for $F_{s}$:
Theorem 5.10.
$\sum_{n=0}^{\infty}\frac{x^{n}}{n!}\left(\sum_{k=0}^{n}F_{s}(n,k)u^{k}\right)=\frac{u^{s}-1}{u-e^{x(u^{s}-1)}}.$
6. Open questions
We have described the game permutation Wordle and its generalization multicolor permutation Wordle. We have also described the strategy CircularShift, which we believe solves both games optimally.
The natural open question is whether CircularShift really is optimal for permutation Wordle, or for multicolor permutation Wordle. Does it dominate all other strategies, in the sense that, for all $r$, it wins more often within $r$ rounds than any other strategy? If not, is it optimal in some more limited sense (e.g., average number of rounds)? Suppose we allow the guesser to guess arbitary strings of (colored) numbers, as in Li–Zhu [LZ24]: does that enable a more efficient strategy?
There is a natural connection between the higher-order $E_{s,z}(n,k)$ and the joint distribution of excedances and fixed points. More precisely, if we let $J_{n}(x,y)=\sum_{\sigma\in S_{n}}x^{\operatorname{exc}(\pi)}y^{\operatorname{fp}(\pi)}$, where $\operatorname{fp}(\pi)$ is the number of fixed points of $\pi$, then
$\sum_{t}E_{s,z}(n,k)t^{k}=s^{n}J_{n}\left(t,\frac{(s-z)t+z}{s}\right).$
Does this joint distribution shed light on these $E_{s,z}$ values, or vice versa?
The authors are indebted to Alex Miller for making the connection with excedances and for proposing the 2-colored variant. We would also like to thank Tim Chow, Keith Frankston, Zachary Hamaker, Ben Howard, Danny Scheinerman, and Jeff VanderKam for helpful discussions.
Appendix A Proving recurrences from generating functions
We now prove Theorem 4.6.
Proof of Theorem 4.6.
When $n=0$, Theorem 4.5 gives
$\sum_{k}D_{()}(0,k)u^{k}=(1-u)\sum_{j}u^{j}=1,$
which matches our initial condition.
For $n>0$, it is convenient to break up the right-hand side of (4) into three pieces:
$X(n,k)$ $:=(k+1){D_{(s_{1},\ldots,s_{n-1})}}(n-1,k)-(k+1-s_{n}){D_{(s_{1},\ldots,s_{n-1})}}(n-1,k-s_{n}),$
$Y(n,k)$ $:=\sum_{r=1}^{s_{n}}{D_{(s_{1},\ldots,s_{n-1})}}(n-1,k-r),$
$Z(n,k)$ $:=\sum_{\ell=1}^{n-1}s_{\ell}D_{{(s_{1},\ldots,\widehat{s_{\ell}},\ldots,s_{n})}}(n-1,k-s_{\ell}).$
We wish to show that, for all $k$, ${D_{(s_{1},\ldots,s_{n})}}(n,k)=X(n,k)+Y(n,k)+Z(n,k)$, or
(8) $\sum_{k}{D_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}=\sum_{k}X(n,k)u^{k}+\sum_{k}Y(n,k)u^{k}+\sum_{k}Z(n,k)u^{k}.$
. We begin with $\sum_{k}X(n,k)u^{k}$. Setting $k^{\prime}=k-s_{n}$,
$\sum_{k}X(n,k)u^{k}$ $=\sum_{k}(k+1){D_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k}-u^{s_{n}}\sum_{k^{\prime}}(k^{\prime}+1){D_{(s_{1},\ldots,s_{n-1})}}(n-1,k^{\prime})u^{k^{\prime}}$
$=(1-u^{s_{n}})\sum_{k}(k+1){D_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k}$
$=(1-u^{s_{n}})\frac{d}{du}\sum_{k}{D_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k+1}$
$=(1-u^{s_{n}})\frac{d}{du}\left[\left(\sum_{j}(j+1)^{n-1}u^{j+1}\right)(1-u)\prod_{\ell=1}^{n-1}(1-u^{s_{\ell}})\right].$
We express this derivative as a sum of three terms:
$\sum_{k}X(n,k)u^{k}=(1-u^{s_{n}})\sum_{j}(j+1)^{n}u^{j}(1-u)\prod_{\ell=1}^{n-1}(1-u^{s_{\ell}})\\ -(1-u^{s_{n}})\sum_{j}(j+1)^{n-1}u^{j+1}\prod_{\ell=1}^{n-1}(1-u^{s_{\ell}})\\ -(1-u^{s_{n}})\sum_{\ell=1}^{n-1}s_{\ell}u^{s_{\ell}-1}\sum_{j}(j+1)^{n-1}u^{j+1}(1-u)\prod_{m\neq\ell,m=1}^{n-1}(1-u^{s_{m}}).$
Each term can now be rewritten using (3):
$\sum_{k}X(n,k)u^{k}=\sum_{k}{D_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}-\frac{u(1-u^{s_{n}})}{1-u}\sum_{k}{D_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k}\\ -\sum_{\ell=1}^{n-1}s_{\ell}u^{s_{\ell}}\sum_{k}D_{(s_{1},\ldots,\widehat{s_{\ell}},\ldots,s_{n})}(n-1,k)u^{k}.$
This is
$\sum_{k}X(n,k)u^{k}=\sum_{k}{D_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}-\sum_{k}Y(n,k)u^{k}-\sum_{k}Z(n,k)u^{k},$
demonstrating (8), as desired. ∎
The proof of Theorem 4.12 uses similar logic.
Proof of Theorem 4.12.
By Theorem 4.11,
$\sum_{k}F_{s}(n,k)u^{k}=(1-u^{s})^{n+1}\sum_{j}(j+1)^{n}u^{j}.$
When $n=0$, this gives $F_{s}(n,k)=(1-u^{s})\sum_{j}u^{j}=1+\cdots+u^{s-1}$, which matches our initial condition. For $n>0$, let
$W(n,k)=(k+1)F_{s}(n-1,k)+(s-k-1)F_{s}(n-1,k-s).$
The calculation
$\sum_{k}W(n,k)u^{k}$ $=\sum_{k}(k+1)F_{s}(n-1,k)u^{k}-u^{s}\sum_{k}(k-s+1)F_{s}(n-1,k-s)u^{k-s}$
$=(1-u^{s})\sum_{k}(k+1)F_{s}(n-1,k)u^{k}$
$=(1-u^{s})\frac{d}{du}u\sum_{k}F_{s}(n-1,k)u^{k}$
$=(1-u^{s})\frac{d}{du}\left[(1-u^{s})^{n}\sum_{j}(j+1)^{n-1}u^{j+1}\right]$
$=(1-u^{s})^{n+1}\sum_{j}(j+1)^{n}u^{j}-nsu^{s-1}(1-u^{s})^{n}\sum_{j}(j+1)^{n-1}u^{j+1}$
$=\sum_{k}F_{s}(n,k)u^{k}-sn\sum_{k}F_{s}(n-1,k)u^{s+k}$
$=\sum_{k}(F_{s}(n,k)-nsF_{s}(n-1,k-s))u^{k},$
completes the proof. ∎