### v3

# Permutation wordle

Abstract.
We introduce a guessing game, “permutation Wordle”, in which a guesser attempts to recover a setter’s hidden permutation of the set $\{1,\ldots,n\}$. In each round, the guesser guesses a permutation, and, as in the game Wordle, learns which entries are correct. We describe a natural strategy, which we conjecture is optimal. We show that the number of permutations this strategy solves in $k+1$ rounds is the Eulerian number $A(n,k)$.
We also describe an extension to multicolor permutations: the setter chooses a permutation and also a coloring of $\{1,\ldots,n\}$. We generalize our strategy, give a recurrence for the number of $s$-colored permutations solved in $k+1$ rounds, and relate these “Wordle numbers” to the Eulerian numbers and to higher-order Eulerian numbers.
1. Introduction
Jon Scieska [9] writes, “You can think of almost everything as a math problem.” We present an example of how this happens. We start with the game Wordle, turn it into a guessing game with permutations, and find connections to a well-studied sequence: the Eulerian numbers.
In 2021, Josh Wardle [12] introduced the online game “Wordle”. In 2022, Wordle was purchased by The New York Times. By 2025, Wordle had been played billions of times.
The goal of Wordle is to discover a secret five-letter word within six guesses. In each round, the guesser chooses a five-letter English word and learns which letters are correct (in the secret word in the same place), which are misplaced (in the secret word in a different place), and which are wrong (not in the secret word at all). [^1] Wordle has well-defined rules for handling repeated letters in the secret word or in the guess, but these are not relevant to this paper.
One can find strategy tips for Wordle online. Optimal play requires an understanding of the set of words in the language. Even partial information about the language helps: for example, typical patterns of consonants and vowels, or common word endings. In order to turn Wordle into a math problem, we make it more abstract.
Our new game is “permutation Wordle”. Choose a whole number $n$. A “word” is a permutation of the set $\{1,\ldots,n\}$: each number appears exactly once. The secret is a permutation, and each guess is also a permutation. The feedback for permutation Wordle is simpler than the feedback for original Wordle: no number is wrong. Each number is either correct or misplaced, because each appears somewhere in the secret.
What is the optimal strategy? How many guesses do we need on average? More generally, what is the probability that we solve after exactly $k$ wrong guesses? In Section 2 we offer a strategy, “CircularShift”, which we conjecture is the best possible. In Section 3 we show that CircularShift solves $A(n,k)$ of the $n!$ permutations in $k+1$ guesses, where $A(n,k)$ are the Eulerian numbers, a sequence that frequently arises in the analysis of permutations [8].
The appearance of Eulerian numbers is a sign that we have found a “natural” math problem from Wordle—it is a satisfying answer. Another sign is that this permutation-guessing problem seems to keep arising. The authors first considered it not because of Wordle, but as an abstraction of permutation-guessing games on Sporcle.com. In July 2025, Richard Stanley [10] posed the same question; he came to it from a guessing game in the app Royal Match.
We can generalize this problem. In Section 4 we introduce a “multicolor” version of the game. We can visualize permutation Wordle as selecting an order for $n$ cards bearing the ranks $1$ through $n$. We generalize to a larger deck of cards, where each rank is present in more than one color; the secret for multicolor permutation Wordle corresponds to an ordered set of $n$ cards where each number appears exactly once. The feedback for the guesser is still a set of which positions are correctly filled, but with a twist: both the number and color must be correct. We adapt CircularShift, and we show in Section 5 that this multicolor problem leads to higher-order Eulerian numbers.
There are connections between Wordle and the classic board game Mastermind; in Section 6 we discuss the existing literature on Mastermind, particularly when the secret is known to be a permutation. Finally, we state some open problems in Section 7.
2. CircularShift
We describe a simple strategy, CircularShift, that solves permutation Wordle in at most $n$ guesses. [^2] Li and Zhu [7] showed that $n$ guesses are necessary. See Section 6 for more discussion. Let $[n]$ denote the set $\{1,\ldots,n\}$, and let $S_{n}$ denote the group of permutations on $[n]$. For now, we act if the setter chooses the secret permutation $\sigma$ uniformly at random. We show in Section 2.2 that we can restrict our analysis to this case, because the guesser can employ a randomized strategy.
We denote the guesser’s guesses by $\{\gamma_{r}\}$ for $r\geq 1$. In round $r$, the feedback
(1) $U_{r}=\{i\mid\gamma_{r}(i)=\sigma(i)\}$
is the set of positions where $\gamma_{r}$ agrees with the setter’s permutation $\sigma.$ Equivalently, $U_{r}$ is the set of fixed points of the permutation $\sigma^{-1}\gamma_{r}$. The guesser may use the information $U_{k}$ for $1\leq k\leq r$ in making the choice of $\gamma_{r+1}$.
2.1. Small $n$
We begin with the case $n=3$. All first guesses are equivalent, so we may as well choose the identity permutation as our first guess $\gamma_{1}$. In response to this first guess, we learn the set of fixed points of $\sigma$.
Figure 1. Permutation Wordle when $n=3$. We solve 11 permutation in one round, 44 in two rounds, and $1$ in three rounds.
Figure 1 shows the different scenarios. There are three possibilities, depending on the number of fixed points of $\sigma.$ If $\sigma$ has three fixed points, then $\sigma$ is the identity and we are done in one guess. If $\sigma$ has exactly one fixed point, then $\sigma$ is one of the three permutations that swaps two values, and the fixed point identifies the swap; in round $2$, we swap those two values and get the correct answer. If $\sigma$ has no fixed points, then $\sigma$ is one of the two $3$-cycles, but we do not know which. Our best strategy is to guess one of the $3$-cycles in round $2$; we will either be correct, or we will need a third guess. Overall, the probability of finding $\sigma$ in one round is $1/6$; in two rounds, $4/6$; in three rounds, $1/6$.
The analysis for $n=4$ is more complicated. Again, let $\gamma_{1}$ be the identity. If $\sigma$ is the identity, we finish in one round. If $\sigma$ has one or two fixed points, we can just keep those values the same in $\gamma_{r}$ for all $r>1$; this lets us reduce to the $n=2$ or $n=3$ case, and we finish in a total of two or three rounds.
What if $\sigma$ is one of the nine derangements, meaning $\sigma$ has no fixed points? One can work out (by hand or by computer exhaustion) that the optimal strategy is to let $\gamma_{2}$ be a $4$-cycle. [^3] This strategy is optimal both in the sense of maximizing the probability of solving within a particular number of rounds and in the sense of minimizing the expected number of rounds. If $\sigma=\gamma_{2}$, we are done in two rounds; if $\sigma^{-1}\gamma_{2}$ has one or two fixed points, we can deduce $\sigma$ in the third round; otherwise, the remaining possibilities for $\sigma$ are $\gamma_{2}^{2}$ and $\gamma_{2}^{3}$, so we choose one of them and solve in either three or four rounds. Figure 2 depicts this overall strategy.
Figure 2. Permutation Wordle when $n=4$. We solve 11 permutation in one round, $11$ in two rounds, $11$ in three rounds, and $1$ in four rounds.
2.2. An adversarial setter
Before we go on to general $n$, we consider the possibility of an adversarial setter. We have described a deterministic strategy, which makes sense if the setter is choosing $\sigma$ uniformly at random. What if the setter uses knowledge of the guesser’s (possibly randomized) strategy to choose the secret permutation? Can the setter make the game harder by choosing $\sigma$ from some other distribution? [^4] We do not consider the case (common in analyses of Mastermind) of the adaptive setter who changes the secret permutation during the game to make the guesser take as long as possible.
For permutation Wordle, the answer turns out to be “no”: we may as well consider a uniform random $\sigma$. We can always randomize our strategy by introducing an extra, uniformly chosen permutation $\tau$. Instead of the sequence of guesses $\gamma_{1}$, $\gamma_{2}$, $\ldots$, we guess $\gamma^{\prime}_{r}=\tau\gamma_{r}$. The feedback in each round is the set of fixed points of $\sigma^{-1}{\gamma^{\prime}_{r}}=(\sigma^{-1}\tau)\gamma_{r}$; this is equivalent to applying a deterministic strategy to $\tau^{-1}\sigma$, which is a uniform random permutation no matter the setter’s strategy.
For the rest of this paper, we consider deterministic strategies, but the reader can imagine that in practice we would multiply by a uniform random $\tau$. The key is that we can analyze average-case behavior. In contrast, in the classic game Mastermind, not all of the setter’s codes are equivalent, so the literature tends to focus on worst-case behavior. See Section 6 for more discussion.
2.3. The general algorithm
We now extend our approach from $n=3$ and $4$ to general $n$. As noted above, we describe a deterministic strategy.
We call our strategy CircularShift. Informally, the idea is as follows: once we get a value correct (that is, $\gamma_{r}(i)=\sigma(i)$), we never change it in subsequent rounds. We take the misplaced values in $\gamma_{r}$ and cycle them around one step to the right to form $\gamma_{r+1}$. Over successive guesses, each value travels to the correct position.
Definition 2.1 (CircularShift).
Let $\gamma_{1}$ be the identity in $S_{n}$. Suppose that, in round $r$, we submit guess $\gamma_{r}$ and receive feedback $U_{r}$ as in (1). If $U_{r}=[n]$, we are done in $r$ rounds. Otherwise, we define $\gamma_{r+1}$ as follows: write $n]\setminus U_{r}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$, and assign
$\gamma_{r+1}(i):=\begin{cases}\gamma_{r}(i)&\text{if $i\in U_{r}$,}\\ \gamma_{r}(a_{j-1})&\text{if $i=a_{j}$ for some $j>1$,}\\ \gamma_{r}(a_{\ell})&\text{if $i=a_{1}$.}\end{cases}$
The approach depicted in Figures 1 and 2 is CircularShift. For a larger example, see Figure 3.
{wordle}
[present color=WordleAbsent]724853169 123456789 821354679 728153469 724853169
Figure 3. The deterministic strategy CircularShift guessing the permutation $\sigma=(7,2,4,8,5,3,1,6,9)$. The strategy succeeds in four rounds.
We conjecture that CircularShift is optimal in the following sense: Suppose that the setter chooses $\sigma\in S_{n}$ uniformly at random. Given a strategy $\mathcal{S}$}, for each $r$, let $P_{r}({\mathcal{S}})$ denote the probability that $\mathcal{S}$} guesses the correct permutation within $r$ guesses.
Conjecture 2.2.
$P_{r}(\textnormal{{CircularShift}})$ attains the maximal value of $P_{r}$ for all $r\geq 1$.
We have verified the conjecture for $n\leq 6$. See Section 7 for further discussion.
Although we cannot prove that CircularShift is optimal, we can completely analyze its cost. To do so, we introduce some numbers that often arise in the combinatorics of permutations.
3. Eulerian numbers
One could write a whole book on the Eulerian numbers [8]. One common definition of the Eulerian numbers is in terms of the number of excedances in a permutation:
Definition 3.1.
For $\pi\in S_{n}$, an excedance is a value $i$ for which $\pi(i)>i$. Let $\operatorname{Exc}(\pi)$ denote the set of excedances: $\operatorname{Exc}(\pi):=\{i\mid\pi(i)>i\}$. Write $\operatorname{exc}(\pi):=\#\operatorname{Exc}(\pi)$, the number of excedances of $\pi$.
Definition 3.2.
The Eulerian number $A(n,k)$ is $\{\pi\in S_{n}\mid\operatorname{exc}(\pi)=k\}$.
For example, when $n=3$, the only permutation with zero excedances is the identity, so $A(3,0)=1$. The only permutation in $S_{3}$ with two excedances is $(2,3,1)$, so $A(3,2)=1.$ We deduce that $A(3,1)=4$.
It is useful to write down a recurrence for the Eulerian numbers.
Proposition 3.3.
The Eulerian numbers $A(n,k)$ satisfy the recurrence
(2) A(n,k)=(k+1)A(n1,k)+(nk)A(n1,k1)A(n,k)=(k+1)A(n-1,k)+(n-k)A(n-1,k-1)
with initial conditions $A(0,0)=1$ and $A(0,k)=0$ for $k\neq 0$.
Proof.
We induct on $n$. The unique element of $S_{0}$ has 0 excedances, verifying the initial condition.
Every permutation $\pi\in S_{n}$ can be uniquely decomposed in the form $\alpha\beta$, where $\beta$ fixes $n$ and $\alpha$ is either the identity, or swaps $n$ with some value $j$. We view $\beta$ as an element of $S_{n-1}$.
How can $\pi$ have exactly $k$ excedances? There are two possibilities. First, $\beta$ could have $k$ excedances, and $\alpha$ is either the identity, or swaps $n$ with one of the $k$ places in $\operatorname{Exc}(\beta)$. Second, $\beta$ could have k1k-1 excedances, and $\alpha$ swaps $n$ with one of the nkn-k places not in $\operatorname{Exc}(\beta)$. This corresponds to the desired recurrence. ∎
The recurrence in Proposition 3.3 implies the reflective symmetry A(n,k)=A(n,nk1)A(n,k)=A(n,n-k-1).
There is also a well-known generating function representation of the Eulerian numbers, which relates to Euler’s work on the subject:
(3) $\frac{\sum_{k}A(n,k)t^{k}}{(1-t)^{n+1}}=\sum_{j}(j+1)^{n}t^{j}.$
One way to derive (3) is to start with the infinite sequence whose enumerator is the right hand side of (3) and then apply the finite difference operator $n+1$ times. The result is a sequence with a polynomial enumerator, that is, a finite number of nonzero elements. As shown for $n=3$ in Figure 4, the nonzero elements of the final sequence are the Eulerian numbers.
[…]
Figure 4. We start with the sequence whose enumerator is $\sum_{j}(j+1)^{n}t^{j}$ (for $n=3$) and apply finite differences (that is, multiply by 1t1-t) four times.
Another common definition of the Eulerian numbers is in terms of descents. A descent of a permutation is a place $i$ with $\pi(i)>\pi(i+1)$; the number of permutations in $S_{n}$ with exactly $k$ descents is also given by $A(n,k)$. One way to prove the equivalence of this alternative formulation for $A(n,k)$ is verify that these numbers satisfy the same recurrence (2). Foata and Schützenberger [3] give an explicit bijection between permutations with $k$ excedances and those with $k$ descents.
The Eulerian numbers show up in many other contexts. For example, let $Y_{1},\ldots,Y_{n}$ be independent uniform random variables in the interval $(0,1)$, and define $X=\left\lfloor{\sum_{i=1}^{n}Y_{i}}\right\rfloor$. The probability that $X=k$ is $A(n,k)/n!$. We now show that these numbers arise in the study of permutation Wordle.
3.1. The main theorem
We will prove the following lemma, which connects CircularShift to excedances:
Lemma 3.4.
The number of guesses it takes CircularShift to guess a permutation $\sigma$ is $1+\operatorname{exc}(\sigma)$.
Lemma 3.4 immediately implies our main theorem:
Theorem 3.5.
The number of permutations solved in $k+1$ guesses by CircularShift is the Eulerian number $A(n,k)$.
Since the Eulerian numbers have reflexive symmetry, we get a simple description of the average-case performance of CircularShift:
Corollary 3.6.
If $\sigma\in S_{n}$ is chosen uniformly at random, then the expected number of rounds for CircularShift is $(n+1)/2$.
This leads us to a weaker version of Conjecture 2.2:
Conjecture 3.7.
For any strategy for permutation Wordle, the average number of rounds is at least $(n+1)/2$.
It remains to prove Lemma 3.4. The idea is to keep track of the sets $R_{k}$ of values placed further to the right by $\gamma_{k}$ than by $\sigma$:
$R_{k}:=\{\gamma_{k}(i):\gamma_{k}(i)>\sigma(i)\}.$
We will show that $R_{k}$ decreases by one each round, starting with $R_{1}=\operatorname{exc}(\sigma)$, until the final round where $R_{1+\operatorname{exc}(\sigma)}=\emptyset$ and $\gamma_{1+\operatorname{exc}(\sigma)}=\sigma$. This is illustrated in Figure 5.
{wordle}
[…]
Figure 5. The deterministic strategy CircularShift guessing the permutation $\sigma=(7,2,4,8,5,3,1,6,9)$. We highlight the values in $R_{k}$ for each $k$; that is, those values that $\gamma_{k}$ places to the right of $\sigma$.
Proof of Lemma 3.4.
Let $\sigma$ be the setter’s permutation. Run CircularShift, producing a series of guesses $\gamma_{1}$, …, $\gamma_{r}$ for some $r$ with $\gamma_{r}=\sigma$. Recall that $U_{k}$ is the set of positions where $\gamma_{k}$ agrees with $\sigma$ and $R_{k}$ is the set of values that $\gamma_{k}$ places further right than $\sigma$ does. (The values in $R_{k}$ are highlighted in Figure 5.) Let $L_{k}$ be the misplaced values that are not in $R_{k}$: the values that $\gamma_{k}$ places further to the left than $\sigma$ does.
Observe that $i\in R_{1}$ if and only if $i>\sigma^{-1}(i)$, meaning $\sigma^{-1}(i)\in\operatorname{Exc}(\sigma)$. We see that $R_{1}=\operatorname{exc}(\sigma).$
Assuming $U_{k}\neq[n]$, how does $R_{k+1}$ differ from $R_{k}$? The set $n]\setminus U_{k}$ has at least two elements, because it is impossible to have exactly one number misplaced. Write $n]\setminus U_{k}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$. Since $\gamma_{k}(a_{\ell})$ is misplaced and it is the rightmost misplaced number, $\gamma_{k}(a_{\ell})\in R_{k}$. Following CircularShift, any values in the correct positions remain in the correct positions: for $i\in U_{k}$, including $1\leq i<a_{1}$, we have $\gamma_{k+1}(i)=\gamma_{k}(i)=\sigma(i)$. Also following CircularShift, the rightmost misplaced value moves to the position of the leftmost misplaced value: $\gamma_{k+1}(a_{1})=\gamma_{k}(a_{\ell})$. So $\gamma_{k}(a_{\ell})\not\in R_{k+1}$.
Any other values in $R_{k}$ move further right, and are also in $R_{k+1}$. Each value in $L_{k}$ moves right, but not beyond its location in $\sigma$, so it is not in $R_{k+1}$. We conclude that $R_{k+1}=R_{k}\setminus\{a_{\ell}\}$, so $R_{k+1}=\#R_{k}-1$.
CircularShift terminates when $U_{r}=[n]$, meaning $\gamma_{r}=\sigma.$ This happens exactly when $R_{r}$ is empty, so the number of rounds for CircularShift is $1+\#R_{1}=1+\operatorname{exc}(\sigma)$. ∎
4. Multicolor Permutation Wordle
We now generalize from the “monochrome” case to a “multicolor” version of the problem. Imagine that, instead of just choosing a permutation of $[n]$, the setter chooses an arrangement of $n$ cards from a larger deck: each card has a rank between $1$ and $n$ and a color. We still require that the setter’s arrangement contain each rank exactly once. More formally:
Definition 4.1.
A multicolor permutation is a permutation in $S_{n}$ together with a coloring $c\colon[n]\to\{0,1,\ldots\}$. An $(s_{1},\ldots,s_{n})$-colored permutation is one where $0\leq c(i)<s_{i}$ for all $i$, and an $s$-colored permutation is one where $0\leq c(i)<s$.
We focus on these cases with explicit bounds on $c(i)$ because the notion of a uniform random $(s_{1},\ldots,s_{n})$-colored or $s$-colored permutation is well defined. We will see that the analysis is roughly the same whether we have $n$ parameters $(s_{1},\ldots,s_{n})$ or a single parameter $s$, but the statements are simpler for a single $s$.
In multicolor permutation Wordle, the setter chooses a multicolor permutation $(\sigma,\kappa)$. The feedback for guess $(\gamma,c)$ is the set of positions $i$ where the guess agrees the setter’s $\sigma$ and $\kappa$, that is, $\gamma(i)=\sigma(i)$ and $c(\gamma(i))=\kappa(\sigma(i))$.
Remark 4.2.
One could reasonably ask whether this is the right rule for the feedback. The honest answer is that this is the rule where the analysis of CircularShift seems to lead to interesting numbers. For example, one interpretation of the Wordle rules would be to get feedback of the form “Red 1 in the right place; blue 2 is misplaced; there is no red 3.” The two-color game with this more informative feedback rule reveals which numbers are red and which blue after just one guess; it is too close to the original “monochrome” permutation Wordle problem.
As in the monochrome version, the argument of Section 2.2 applies: we may assume that the setter chooses $\sigma$ and $\kappa$ uniformly at random, and we may employ a deterministic strategy. Our first guess is $(\gamma_{1},c_{1})$ where $\gamma_{1}$ is the identity permutation and $c_{1}(i)=0$ for all $i$.
4.1. CircularShift for multicolor permutation Wordle
We now generalize from monochrome CircularShift to the multicolor setting. We rotate the values of $\gamma_{r}$ just as before, and we start with all colors equal to 0. When a value with color $a$ loops all the way around to its starting position, that means the color is incorrect, so we increment the color to $a+1$ as the value $i$ begins its next circuit.
Definition 4.3 (CircularShift for multicolor permutation Wordle).
Let $\gamma_{1}$ be the identity in $S_{n}$, and let $c_{1}\equiv 0$. Suppose that, in round $r$, we submit guess $(\gamma_{r},c_{r})$ and receive feedback
$U_{r}=\{i\mid\gamma_{r}(i)=\sigma_{r}(i)\ \mathrm{and}\ c_{r}(\gamma_{r}(i))=\kappa(\sigma(i))\},$
the set of positions $i$ where $(\gamma_{r},c_{r})$ and $(\sigma,\kappa)$ agree on both the value and the color of that value. If $U_{r}=[n]$, we are done in $r$ rounds. Otherwise, we define $\gamma_{r+1}$ and $c_{r+1}$ as follows: write $n]\setminus U_{r}=\{a_{1},\ldots,a_{\ell}\}$ with $a_{j}<a_{j+1}$. Let
$\gamma_{r+1}(i):=\begin{cases}\gamma_{r}(i)&\text{if $i\in U_{r}$,}\\ \gamma_{r}(a_{j-1})&\text{if $i=a_{j}$ for some $j>1$,}\\ \gamma_{r}(a_{\ell})&\text{if $i=a_{1}$.}\end{cases}$
Also,
[…]
We next prove a multicolor analogue of Lemma 3.4.
Lemma 4.4.
The number of guesses it takes CircularShift to guess a multicolor permutation $(\sigma,\kappa)$ is
$1+\operatorname{exc}(\sigma)+\sum_{i=1}^{n}\kappa(i).$
The idea of the proof is to count down the number of rounds remaining. We introduce quantities $Z_{k,i}$ representing the number of times after round $k$ that the value $i$ will wrap around before we finish solving. In each round, one value ($i=\gamma_{k}(a_{\ell})$) wraps around, so the total $\sum_{i}Z_{k,i}$ decreases by $1$. We are done when the countdown reaches zero. See Figure 6 for an example.
{wordle}
[…]
Figure 6. Applying CircularShift to two-color permutation Wordle with $n=5$. Here color 0 is shown in white, and color 11 in black. The highlights show how often each number will need to wrap around; note that the 55 is highlighted twice, once because of its location and once because of its color.
Proof.
As in the proof of Lemma 3.4, we construct a quantity which decreases by $1$ in each round and which is 0 when we have the right answer. Suppose that, in round $k$, we have some guess $(\gamma_{k},c_{k})$. For each $i$, let $\delta(i)=1$ when $i$ occurs further to left in $\sigma$ than in the identity; that is, $i>\sigma^{-1}(i)$, or $\sigma^{-1}(i)$ is an excedance of $\sigma$. Let $\epsilon_{k}(i)=1$ when $i>\gamma_{k}^{-1}(i)$. Let
$Z_{k,i}=\kappa(i)-c_{k}(i)+\delta(i)-\epsilon_{r}(i)$
and let $Z_{k}=\sum_{i=1}^{n}Z_{k,i}$. We show that $Z_{1}=\operatorname{exc}(\sigma)+\sum_{i}\kappa(i),$ that $(\gamma_{k},c_{k})\neq(\sigma,\kappa)$ implies $Z_{k}>0,$ and that $Z_{k+1}=Z_{k}-1$, as illustrated by Figure 6.
Since $\gamma_{1}$ is the identity and $c_{1}\equiv 0$, we see that $Z_{1,i}\geq 0$ for all values $i$ and $Z_{1}=\operatorname{exc}(\sigma)+\sum_{i}\kappa(i)$. Note also that $\kappa(i)\geq c_{k}(i)$ for all $k$ and $i$, because, according to CircularShift, $c_{k}(i)$ is a nondecreasing sequence in $k$, and $c_{k+1}(i)=c_{k}(i)+1$ implies that value $i$ with color $c_{k}(i)$ visited every position in $n]\setminus U_{k}$ in earlier rounds. When $\gamma_{k}^{-1}(i)\in U_{k},$, we have $\delta(i)=\epsilon_{k}(i)$ and $\kappa(i)=c_{k}(i)$, so $Z_{k,i}=Z_{k+1,i}=0.$ When $U_{k}=[n]$, $\gamma_{k}=\sigma$, $c_{k}=\kappa$, and $Z_{k}=0$.
Suppose $U_{k}\subsetneq[n].$ Enumerate $n]\setminus U_{k}$ as an increasing sequence $a_{1},\ldots,a_{\ell}$. How do $Z_{k,i}$ and $Z_{k+1,i}$ compare when $i=\gamma_{k}(a_{j})$ for $j<\ell$? If $a_{j}<i\leq a_{j+1}$, then $c_{k+1}(i)=c_{k}(i)+1$, $\epsilon_{k}(i)=0$, and $\epsilon_{k+1}(i)=1$. Otherwise, $c_{k+1}(i)=c_{k}(i)$ and $\epsilon_{k+1}(i)=\epsilon_{k}(i)$. Either way, $Z_{k+1,i}=Z_{k,i}.$
It remains to consider $Z_{k,i}$ and compare $Z_{k+1,i}$ and $Z_{k,i}$ for $i=\gamma_{k}(a_{\ell})$. There are three possibilities. If $i>a_{\ell}\geq a_{1}$, then $\epsilon_{k+1}(i)=\epsilon_{k}(i)=1$ and $c_{k+1}(i)=c_{k}(i)+1$. If $a_{\ell}\geq a_{1}\geq i$, then $\epsilon_{k+1}(i)=\epsilon_{k}(i)=0$ and $c_{k+1}(i)=c_{k}(i)+1$. If $a_{\ell}\geq i>a_{1}$, then $\epsilon_{k+1}(i)=1,$ $\epsilon_{k}(i)=0$, $c_{k+1}(i)=c_{k}(i)$, and $c_{k}(i)<\kappa(i)+\delta(i).$ In each case, $Z_{k,i}>0$ and $Z_{k+1,i}=Z_{k,i}-1$. Since this is the only $i$ for which $Z_{k+1,i}\neq Z_{k,i}$, we have $Z_{k+1}=Z_{k}-1$. The positive value of $Z_{k,i}$ verifies that $(\gamma_{k},c_{k})\neq(\sigma,\kappa)$ implies $Z_{k}>0.$
At round $r=1+\operatorname{exc}(\sigma)+\sum_{i}\kappa(i)$, we have $Z_{r}=0$. and, consequently, $(\gamma_{r},c_{r})=(\sigma,\kappa)$.
Conjecture 4.5.
For each $r$, the probability that CircularShift solves a uniformly chosen $(s_{1},\ldots,s_{n})$-colored permutation within $r$ rounds is optimal, that is, it is at least as large as for any other strategy.
4.2. Multicolor recurrences and generating functions
To understand the expected behavior of CircularShift on multicolor permutations with various numbers of colors, it is convenient to work with generating functions. On Page 3, we mentioned the standard representation of the Eulerian numbers (3), restated here:
$\frac{\sum_{k}A(n,k)t^{k}}{(1-t)^{n+1}}=\sum_{j}(j+1)^{n}t^{j}.$
We generalize this identity to the “Wordle numbers”. We prove our results in the general multicolor form, and then give the simpler expressions we get in the $s$-colored case.
Definition 4.6.
We let ${W_{(s_{1},\ldots,s_{n})}}(n,k)$ denote the number of $(s_{1},\ldots,s_{n})$-colored permutations of length $n$ for which CircularShift succeeds in exactly $k+1$ rounds. We let $W_{s}(n,r)$ denote the number of $s$-colored permutations for which CircularShift succeeds in exactly $r+1$ rounds.
Theorem 4.7.
(4) $\frac{\sum_{k}{W_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}}{(1-u)\prod_{\ell}(1-u^{s_{\ell}})}=\sum_{j}(j+1)^{n}u^{j}.$
Proof.
By Lemma 4.4,
$\sum_{k}{W_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}$ $=\sum_{\sigma,\kappa}u^{\operatorname{exc}(\sigma)+\kappa(1)+\cdots+\kappa(n)}$
$=\left(\sum_{\sigma}u^{\operatorname{exc}(\sigma)}\right)\left(\sum_{\kappa}\prod_{\ell=1}^{n}u^{\kappa(\ell)}\right)$
$=\left(\sum_{k}A(n,k)u^{k}\right)\prod_{\ell=1}^{n}(1+u+\cdots+u^{s_{\ell}-1}),$
since $A(n,k)$ counts excedances. Using (3),
$\sum_{k}{W_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}$ $=\left(\sum_{j}(j+1)^{n}u^{j}\right)(1-u)^{n+1}\prod_{\ell=1}^{n}\frac{1-u^{s_{\ell}}}{1-u}.$
The result follows. ∎
From the generating function (4) we can derive a recurrence for ${W_{(s_{1},\ldots,s_{n})}}(n,k)$. We use $(s_{1},\ldots,\widehat{s_{\ell}},\ldots,s_{n})$} to denote the (n1)(n-1)-long sequence obtained from $(s_{1},\ldots,s_{n})$ by removing $s_{\ell}$.
Theorem 4.8.
The counts ${W_{(s_{1},\ldots,s_{n})}}(n,k)$ satisfy the recurrence
[…]
with initial conditions $W_{()}(0,0)=1$ and $W_{()}(0,k)=0$ for $k\neq 0$.
The proof uses standard generating function techniques; see Appendix A.
Theorems 4.7 and 4.8 are both simpler in the $s$-colored case:
Theorem 4.9.
$\frac{\sum_{k}W_{s}(n,k)u^{k}}{(1-u)(1-u^{s})^{n}}=\sum_{j}(j+1)^{n}u^{j}.$
Theorem 4.10.
The counts $W_{s}(n,k)$ satisfy the recurrence
(6) $W_{s}(n,k)=(k+1)W_{s}(n-1,k)+(sn-k)W_{s}(n-1,k-s)+\sum_{\ell=1}^{s-1}W_{s}(n-1,k-\ell)$
with initial conditions $W_{s}(0,0)=1$ and $W_{s}(0,k)=0$ for $k\neq 0$.
Note that, when $s=1$, this expression  (6) is the recurrence (2) for the Eulerian numbers.
The recurrence (6) has a natural symmetry:
Corollary 4.11.
We have $W_{s}(n,sn-r-1)=W_{s}(n,r)$.
4.3. Sums of consecutive values
We can find patterns in the sums of consecutive values of these sequences $W_{s}(n,r)$. For example, consider adding pairs of numbers from $W_{2}(n,r)$; the case $n=4$ is shown in Figure 7. If we start with the first two numbers, then the next two, and so on, we obtain multiples of the Eulerian numbers: the sequence is $2^{n}A(n,r)$. However, if we instead start with just one number, then two at a time, we obtain a different known combinatorial sequence: the Eulerian numbers of type B [1]. We now develop a recurrence for these sums of consecutive values of Wordle numbers; in Section 5, we will show that these sums give the “higher-order Eulerian numbers” [11].
Remark 4.12.
Given two sequences with the same sum, it is not surprising that one can write down a sequence that simultaneously refines both of them; that is, the numbers can add to one or the other depending on how we pair them up. What is surprising is that this particular sequence $W_{2}(n,r)$ arises in its own right, in the analysis of two-color CircularShift. One is naturally led to wonder what else these $W_{s}$ numbers might be counting.
11561115115611511617617616176230761
Figure 7. Two ways to group the sequence $W_{2}(4,k)$.
Definition 4.13.
For each $s,n,k$, we define $F_{s}(n,k)$ to be the sum of $s$ consecutive values of $W_{s}(n,r)$:
$F_{s}(n,k)=\sum_{\ell=0}^{s-1}W_{s}(n,k-\ell).$
Theorem 4.14.
(7) $\frac{\sum_{k}F_{s}(n,k)u^{k}}{(1-u^{s})^{n+1}}=\sum_{j}(j+1)^{n}u^{j}.$
Proof.
From the definition, $\sum_{k}F_{s}(n,k)u^{k}=(1+\cdots+u^{s-1})\sum_{k}W_{s}(n,k)u^{k}$. The result follows immediately from Theorem 4.9. ∎
Theorem 4.15.
The numbers $F_{s}(n,k)$ satisfy the recurrence
$F_{s}(n,k)=(k+1)F_{s}(n-1,k)+(s(n+1)-k-1)F_{s}(n-1,k-s).$
with initial conditions $F_{s}(0,k)=1$ when $0\leq k<s$ and $F_{s}(0,k)=0$ otherwise.
Again, a proof can be found in Appendix A.
Remark 4.16.
The sequences ${W_{(s_{1},\ldots,s_{n})}}(n,\cdot)$ and $F_{s}(n,\cdot)$ are log-concave. Note that ${W_{(s_{1},\ldots,s_{n})}}(n,\cdot)$ is the iterated convolution of $A(n,\cdot)$ with sequences $(1,\ldots,1)$ of length $s_{\ell}$ for $1\leq\ell\leq n$, and $F_{s}(n,\cdot)$ is the convolution of $W_{s}(n,\cdot)$ with $(1,\ldots,1)$. The convolution of two log-concave sequences is log-concave [5].
5. Higher-order Eulerian numbers
5.1. Multiples of Eulerian numbers
As noted above, sums of consecutive values of $W_{2}(n,r)$ recover the values $2^{n}A(n,r)$. We begin our discussion of higher-order Eulerian numbers by generalizing this observation to arbitrary $s$.
Proposition 5.1.
For all $n$ and $k$,
$F_{s}(n,sk+s-1)=s^{n}A(n,k).$
Proof.
For $n=0$, $F_{s}(n,sk+s-1)=1$ when $k=0$ and $F_{s}(n,sk+s-1)=0$ otherwise.
For $n>0$, by Theorem 4.15,
$F_{s}(n,sk+s-1)$ $=(sk+s)F_{s}(n-1,sk+s-1)+(s(n+1)-sk-s)F_{s}(n-1,sk-1)$
$=s\left[(k+1)F_{s}(n-1,sk+s-1)+(n-k)F_{s}(n-1,s(k-1)+s-1)\right].$
By Proposition 3.3, this same recurrence is satisfied by $s^{n}A(n,k)$. ∎
Corollary 5.2.
The probability that CircularShift succeeds on $s$-colored permutation Wordle in at most $sr$ rounds is the same as the probability that CircularShift succeeds on permutation Wordle in at most $r$ rounds.
5.2. $z$-excedances
To describe the other possible sums of $W_{s}(n,r)$, we need to introduce the higher-order Eulerian numbers. We follow (and slightly generalize) the approach of Steingrímsson [11]. Let $S_{n,s}$ denote the set of $s$-colored permutations of length $n$.
Definition 5.3.
For $(\pi,c)\in S_{n,s}$, a $z$-excedance is a value $j$ for which either $\pi(j)>j$, or for which $\pi(j)=j$ and $c(\pi(j))\geq z$. Equivalently, it is a value $j$ where $(\pi(j),c(\pi(j)))\succ(i,z-1)$ lexicographically.
For $1\leq z\leq s$, we let $E_{s,z}(n,k)$ denote the number of $s$-colored permutations of length $n$ with $k$ $z$-excedances.
Steingrímsson analyzes the case $z=1$ and derives the recurrence
$E_{s,1}(n,k)=(sk+1)E_{s,1}(n-1,k)+(s(n-k)+s-1)E_{s,1}(n-1,k-1)$
and the generating function
(8) $\frac{\sum_{k}E_{s,1}(n,k)t^{k}}{(1-t)^{n+1}}=\sum(sj+1)^{n}t^{j}.$
For $s=2$, these are the Eulerian numbers of type B.
We generalize Steingrímsson’s results to larger values of $z$:
Theorem 5.4.
The numbers $E_{s,z}(n,k)$ satisfy the recurrence
$E_{s,z}(n,k)=(sk+z)E_{s,z}(n-1,k)+(s(n-k)+s-z)E_{s,z}(n-1,k-1)$
with initial conditions $E_{s,z}(0,k)=1$ when $k=0$ and $E_{s,z}(0,k)=0$ for $k\neq 0$.
The proof follows the same lines as that of Proposition 3.3.
Proof.
We induct on $n$; the unique element of $S_{0}$ has 0 $z$-excedances, verifying the initial condition.
Every $s$-colored permutation $(\pi,c)\in S_{n,s}$ can be uniquely constructed as follows: start with an $s$-colored permutation $(\beta,b)\in S_{n-1,s}$. Extend $b$ to $c$ by $c(i)=b(i)$ for $i<n$ and making a choice $c(n)$ of a color for value $n$. To make $\pi$, extend $\beta$ to $S_{n}$ by making $n$ a fixed point, then let $\pi=\alpha\beta$ with $\alpha$ either the identity or the swap of $i$ and $n$ for some $i<n$.
How can $(\pi,c)$ have exactly $k$ $z$-excedances? There are two possibilities. First, $(\beta,b)$ could have $k$ $z$-excedances, and either $\alpha$ swaps $n$ with one of the $k$ $z$-excedances (in which case any color $c(n)$ works), or $\alpha$ leaves $n$ in place and $c(n)<z$. Given $(\beta,b)$ with $k$ $z$-excedances, this yields $sk+z$ $s$-colored permutations $(\pi,c)$.
Second, $(\beta,b)$ could have k1k-1 $z$-excedances, and either $\alpha$ swaps $n$ with one of the nkn-k non-$z$-excedances (in which case any color $c(n)$ works), or $\alpha$ leaves $n$ in place and $c(n)\geq z$ (giving one more $z$-excedance). Given $(\beta,b)\in S_{n-1,s}$ with k1k-1 $z$-excedances, this yields s(nk)+szs(n-k)+s-z $s$-colored permutations $(\pi,c)\in S_{n,s}$.
The sum of these two cases yields the desired recurrence. ∎
Note that $E_{s,s}(n,k)$ is simply $s^{n}A(n,k)$; since no color can exceed s1s-1, the colors play no role and we simply count excedances. When $z=s$, the recurrence of Theorem 5.4 simplifies to
$E_{s,s}(n,k)=s(k+1)E_{s,s}(n-1,k)+s(n-k)E_{s,s}(n-1,k-1),$
which is the recurrence for $F_{s}(n,sk+s-1)$ discussed in Proposition 5.1. We can apply similar logic to any value of $z$.
Proposition 5.5.
For $1\leq z\leq s$, we have
$F_{s}(n,sk+z-1)=E_{s,z}(n,k).$
Proof.
When $n=0$, $F_{s}(n,sk+z-1)$ is $1$ when $k=0$ and 0 otherwise, as desired.
For $n>0$, by Theorem 4.15,
$F_{s}(n,sk+z-1)=(sk+z)F_{s}(n-1,sk+z-1)+(s(n-k)+s-z)F_{s}(n-1,s(k-1)+z-1).$
By Proposition 5.4, this is the recurrence satisfied by $E_{s,z}(n,k)$. ∎
Proposition 5.6.
$E_{ds,dz}(n,k)=d^{n}E_{s,z}(n,k).$
Proof.
The proof is the same as that of Proposition 5.1, except that we pull out a factor of $d$, rather than $s$, each time we apply the recurrence. ∎
Remark 5.7.
Since $E_{s,z}(n,\cdot)$ is a subsequence of the log-concave $F(n,\cdot)$ (see Remark 4.16), $E_{s,z}(n,\cdot)$ is also log-concave.
Theorem 5.8.
For $1\leq z\leq s$,
$\frac{\sum_{k}E_{s,z}(n,k)t^{k}}{(1-t)^{n+1}}=\sum_{j}(sj+z)^{n}t^{j}.$
Note that, when $z=1$, Theorem 5.8 recovers Steingrímsson’s (8).
Proof.
By Theorem 4.14,
$\sum_{k}F_{s}(n,k)u^{k}=(1-u^{s})^{n+1}\sum_{j}(j+1)^{n}u^{j}.$
On each side, consider only terms corresponding to $u^{k}$ with $k\equiv z-1\pmod{s}$:
$\sum_{k}F_{s}(n,sk+z-1)u^{sk+z-1}=(1-u^{s})^{n+1}\sum_{j}(sj+z)^{n}u^{sj+z-1}.$
We divide both sides by $u^{z-1}$ and set $t=u^{s}$: with Proposition 5.5, this proves the identity. ∎
We can also prove a version of the classic identity
$\sum_{n=0}^{\infty}\frac{x^{n}}{n!}\left(\sum_{k=0}^{n}A(n,k)t^{k}\right)=\frac{t-1}{t-e^{x(t-1)}}$
and of Steíngrimsson’s similar identity for $E_{s,1}(n,k)$ [11].
Corollary 5.9.
$\sum_{n=0}^{\infty}\frac{x^{n}}{n!}\left(\sum_{k=0}^{n}E_{s,z}(n,k)t^{k}\right)=\frac{(1-t)e^{zx(1-t)}}{1-te^{sx(1-t)}}.$
Using Theorem 4.14, we can do the same for $F_{s}$:
Theorem 5.10.
$\sum_{n=0}^{\infty}\frac{x^{n}}{n!}\left(\sum_{k=0}^{n}F_{s}(n,k)u^{k}\right)=\frac{u^{s}-1}{u-e^{x(u^{s}-1)}}.$
6. Connections with Mastermind
There are natural connections between Wordle and the classic game Mastermind, invented by Mordecai Meirowitz. In Mastermind, the setter chooses a “code”: a string of length $n$ from an alphabet of size $m$ (which we can write as $[m]$). In each round, the guesser tries a string in $m]^{n}$, and the setter responds with two numbers: the number of entries that are correct (right number, right place) and the number of additional entries that are misplaced (right number, wrong place). This is a similar setup to Wordle, except that in Wordle we find out specifically which entries are correct or misplaced, instead of just a count.
Donald Knuth [6] showed that, for the original game ($n=4$ and $m=6$), five rounds are sufficient for the guesser to guess the code. Many others have evaluated the game for various values of $n$ and $m$. The typical emphasis is on a worst-case analysis; one can imagine the setter adversarially altering the code to be consistent with all information presented thus far, with the goal of maximizing the number of rounds. One reason for this worst-case analysis is that not all codes are the same; even if we do something along the lines of Section 2.2, permuting the colors and the entries, there is a difference between the code $(1,2,3,4)$ (with no repeated entries) and $(1,1,2,3)$ (with one repeat). An “average-case” analysis only makes sense with respect to some distribution on the setter’s choices.
One Mastermind variant is “permutation Mastermind”, in which $m=n$ and both the setter’s code and the guesser’s guesses must contain each value exactly once. (In this variant, the setter responds with the number $c$ of correct entries, since the number misplaced is simply ncn-c.) El Ouali et al. [2] have shown that $n\log_{2}n+O(n)$ queries suffice.
Li and Zhu [7] have analyzed a generalized version of Wordle, with a code from the alphabet $m]^{n}$; essentially, they are playing Mastermind with Wordle-like feedback. Among other things, they show that, in the permutation setting where $m=n$ and the code is required to use each value exactly once, then the guesser needs $n$ guesses in the worst case; this is true even if the guesses are are allowed to have values multiple times. This worst-case bound applies directly to our setting; their result shows that, for any permutation Wordle strategy, there must be some input where it requires $n$ guesses.
Li and Zhu give a simple strategy with a matching upper bound of $n$ guesses for the general case $m=n$: the first n1n-1 guesses are the strings $1^{n}$, $2^{n}$, $\ldots$, $(n-1)^{n}$, at which point the guesser knows the code. Our result can be considered an improvement on this simple strategy, in that it has the same worst-case behavior but better average-case behavior.
7. Open questions
We have described the game permutation Wordle and its generalization multicolor permutation Wordle. We have also described the strategy CircularShift, which we believe solves both games optimally.
The natural open question is whether CircularShift really is optimal for permutation Wordle, or for multicolor permutation Wordle. Does Conjecture 2.2 hold—that is, does CircularShift dominate all other strategies, in the sense that, for all $r$, it wins more often within $r$ rounds than any other strategy? If not, is it optimal in some more limited sense (e.g., average number of rounds)? Suppose we allow the guesser to guess arbitrary strings of (multicolor) numbers, as in Li–Zhu [7]: does that enable a more efficient strategy?
Aurora Hiveley [4] has done some computer analyses of CircularShift, and also proven Conjecture 2.2 in the case $r\leq 3$. It would be interesting to see how far her work can be extended.
We introduced the Wordle numbers $W_{s}(n,r)$, and showed that they simultaneously refine the Eulerian numbers $A(n,k)$ and the higher-order $E_{s,z}(n,k)$. Do these numbers count anything on their own, outside of the permutation Wordle setting?
There is a natural connection between the higher-order $E_{s,z}(n,k)$ and the joint distribution of excedances and fixed points. More precisely, if we let $J_{n}(x,y)=\sum_{\sigma\in S_{n}}x^{\operatorname{exc}(\pi)}y^{\operatorname{fp}(\pi)}$, where $\operatorname{fp}(\pi)$ is the number of fixed points of $\pi$, then
$\sum_{t}E_{s,z}(n,k)t^{k}=s^{n}J_{n}\left(t,\frac{(s-z)t+z}{s}\right).$
Does this joint distribution shed light on these $E_{s,z}$ values?
Acknowledgments
The authors are indebted to Alex Miller for making the connection with excedances and for proposing the two-colored variant. We would also like to thank Joe Buhler, Tim Chow, Keith Frankston, Zachary Hamaker, Ben Howard, Danny Scheinerman, Jeff VanderKam, and Doron Zeilberger for helpful discussions. And we appreciate the package wordle.sty, by Andrew Mathas and Cédric Pierquet, which we used to make the Wordle diagrams throughout the paper.
Appendix A Proving recurrences from generating functions
We now prove Theorem 4.8.
Proof of Theorem 4.8.
When $n=0$, Theorem 4.7 gives
$\sum_{k}W_{()}(0,k)u^{k}=(1-u)\sum_{j}u^{j}=1,$
which matches our initial condition.
For $n>0$, it is convenient to break up the right-hand side of (5) into three pieces:
$X(n,k)$ $:=(k+1){W_{(s_{1},\ldots,s_{n-1})}}(n-1,k)-(k+1-s_{n}){W_{(s_{1},\ldots,s_{n-1})}}(n-1,k-s_{n}),$
$Y(n,k)$ $:=\sum_{r=1}^{s_{n}}{W_{(s_{1},\ldots,s_{n-1})}}(n-1,k-r),$
$Z(n,k)$ $:=\sum_{\ell=1}^{n-1}s_{\ell}W_{{(s_{1},\ldots,\widehat{s_{\ell}},\ldots,s_{n})}}(n-1,k-s_{\ell}).$
We wish to show that, for all $k$, ${W_{(s_{1},\ldots,s_{n})}}(n,k)=X(n,k)+Y(n,k)+Z(n,k)$, or
(9) $\sum_{k}{W_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}=\sum_{k}X(n,k)u^{k}+\sum_{k}Y(n,k)u^{k}+\sum_{k}Z(n,k)u^{k}.$
. We begin with $\sum_{k}X(n,k)u^{k}$. Setting $k^{\prime}=k-s_{n}$,
$\sum_{k}X(n,k)u^{k}$ $=\sum_{k}(k+1){W_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k}-u^{s_{n}}\sum_{k^{\prime}}(k^{\prime}+1){W_{(s_{1},\ldots,s_{n-1})}}(n-1,k^{\prime})u^{k^{\prime}}$
$=(1-u^{s_{n}})\sum_{k}(k+1){W_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k}$
$=(1-u^{s_{n}})\frac{d}{du}\sum_{k}{W_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k+1}$
$=(1-u^{s_{n}})\frac{d}{du}\left[\left(\sum_{j}(j+1)^{n-1}u^{j+1}\right)(1-u)\prod_{\ell=1}^{n-1}(1-u^{s_{\ell}})\right].$
We express this derivative as a sum of three terms:
$\sum_{k}X(n,k)u^{k}=(1-u^{s_{n}})\sum_{j}(j+1)^{n}u^{j}(1-u)\prod_{\ell=1}^{n-1}(1-u^{s_{\ell}})\\ -(1-u^{s_{n}})\sum_{j}(j+1)^{n-1}u^{j+1}\prod_{\ell=1}^{n-1}(1-u^{s_{\ell}})\\ -(1-u^{s_{n}})\sum_{\ell=1}^{n-1}s_{\ell}u^{s_{\ell}-1}\sum_{j}(j+1)^{n-1}u^{j+1}(1-u)\prod_{m\neq\ell,m=1}^{n-1}(1-u^{s_{m}}).$
Each term can now be rewritten using (4):
$\sum_{k}X(n,k)u^{k}=\sum_{k}{W_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}-\frac{u(1-u^{s_{n}})}{1-u}\sum_{k}{W_{(s_{1},\ldots,s_{n-1})}}(n-1,k)u^{k}\\ -\sum_{\ell=1}^{n-1}s_{\ell}u^{s_{\ell}}\sum_{k}W_{(s_{1},\ldots,\widehat{s_{\ell}},\ldots,s_{n})}(n-1,k)u^{k}.$
This is
$\sum_{k}X(n,k)u^{k}=\sum_{k}{W_{(s_{1},\ldots,s_{n})}}(n,k)u^{k}-\sum_{k}Y(n,k)u^{k}-\sum_{k}Z(n,k)u^{k},$
demonstrating (9), as desired. ∎
The proof of Theorem 4.15 uses similar logic.
Proof of Theorem 4.15.
By Theorem 4.14,
$\sum_{k}F_{s}(n,k)u^{k}=(1-u^{s})^{n+1}\sum_{j}(j+1)^{n}u^{j}.$
When $n=0$, this gives $\sum F_{s}(0,k)u^{k}=(1-u^{s})\sum_{j}u^{j}=1+\cdots+u^{s-1}$, which matches our initial condition $F_{s}(0,k)=1$ for $0\leq k<s$ and $F_{s}(0,k)=0$ otherwise. For $n>0$, let
$V(n,k)=(k+1)F_{s}(n-1,k)+(s-k-1)F_{s}(n-1,k-s).$
The calculation
$\sum_{k}V(n,k)u^{k}$ $=\sum_{k}(k+1)F_{s}(n-1,k)u^{k}-u^{s}\sum_{k}(k-s+1)F_{s}(n-1,k-s)u^{k-s}$
$=(1-u^{s})\sum_{k}(k+1)F_{s}(n-1,k)u^{k}$
$=(1-u^{s})\frac{d}{du}u\sum_{k}F_{s}(n-1,k)u^{k}$
$=(1-u^{s})\frac{d}{du}\left[(1-u^{s})^{n}\sum_{j}(j+1)^{n-1}u^{j+1}\right]$
$=(1-u^{s})^{n+1}\sum_{j}(j+1)^{n}u^{j}-nsu^{s-1}(1-u^{s})^{n}\sum_{j}(j+1)^{n-1}u^{j+1}$
$=\sum_{k}F_{s}(n,k)u^{k}-sn\sum_{k}F_{s}(n-1,k)u^{s+k}$
$=\sum_{k}(F_{s}(n,k)-nsF_{s}(n-1,k-s))u^{k},$
completes the proof. ∎